### v1

# Higher-Order Staircase Codes

Abstract
We generalize staircase codes and tiled diagonal zipper codes, preserving their key properties while allowing each coded symbol to be protected by arbitrarily many component codewords rather than only two. This generalization arises from the marriage of two distinct combinatorial objects: difference triangle sets and finite-geometric nets, which have typically been applied separately to code design. We demonstrate one realization of the proposed codes, obtaining powerful, high-rate, low-error-floor, and low-complexity coding schemes based on simple iterative syndrome-domain decoding of coupled Hamming component codes. We anticipate that the proposed codes improve performance–complexity–latency tradeoffs in high-throughput communications applications, most notably fiber-optic, in which classical staircase codes and zipper codes have been applied.
Index Terms:
Staircase codes, product codes, finite-geometric nets, difference triangle sets, iterative algebraic decoding.
I Overview
Staircase codes [2] and related product-like codes [3] have provided a highly-competitive paradigm for energy-efficient, high-rate, high-throughput, and high-reliability, error control coding. Such codes use syndrome-domain iterative algebraic or bounded-distance decoding of algebraic component codes to enable high decoder throughputs, while maintaining relatively low internal decoder data-flow. Furthermore, their block-convolutional structures translate to natural pipelining and parallelism, the combination of which is essential for efficient hardware designs. Moreover, their typically combinatorial constructions ensure control over error floors which must usually be below $10^{-15}$ in the relevant applications. Most notably, such codes are used in fiber-optic communication [4, Ch. 7], but are also relevant to flash storage applications [5] which increasingly demand similar code characteristics.
In this paper, we provide a fruitful generalization of staircase codes which arises from the combination of difference triangle sets (DTSs) [6, Part VI: Ch. 19] and finite-geometric nets [6, Part III: Ch. 3]. Both DTSs and nets are well-studied combinatorial objects found in the standard reference [6] and both have been separately applied to code design, e.g., in [7, 8] and [9, 10] respectively. In these examples, DTSs lead naturally to convolutional codes while nets lead naturally to block codes. It is unsurprising then that their combination leads to a generalization of staircase codes which themselves combine aspects of block and convolutional coding.
We will proceed to summarize, for the reader’s convenience, the key construction and observations of this paper. A higher-order staircase code is determined by three intertwined objects: a DTS, a net, and a component code. Each of these objects is characterized by two parameters, the meanings of which will be defined later. The relationships between these parameters and properties of the resulting code are as follows: Given an $(L,M)$-DTS, an $(M+1,S/L)$-net, and a component code of length $(M+1)S$ and dimension $(M+1)S-r$, one obtains a rate $1-r/S$ spatially-coupled code in which every symbol is protected by $M+1$ component codewords and any pair of distinct component codewords share at most one symbol. Alternatively, the Tanner graph of the code is an infinite semiregular $4$-cycle-free bipartite graph with variable node degree $M+1$ and generalized constraint node degree $(M+1)S$. The relevance of the parameter $L$ is that increasing it can reduce the encoding and decoding memory at the cost of reduced parallelism while keeping all of the aforementioned code properties completely unchanged.
The proposed code family recovers as particular cases some existing codes and interpolates between them:
  • When $L=M=1$, the classical staircase codes of [2] are recovered.
  • When $L>1$ and $M=1$, the tiled diagonal zipper codes of [3] are recovered.
  • When $S=L$ and $r=1$, we recover a recursive (rather than feedforward) version of the self-orthogonal convolutional codes of [7].
When $M>1$ and $L=1$, we obtain a generalization of staircase codes with higher variable degree which was included in a conference version of this work [1]. A similar construction is included in the recently issued US patent [11], which recognizes the role of DTSs in such a generalization of staircase codes but makes no explicit use of finite-geometric nets. Importantly, when $L>1$ and $M>1$, our construction becomes a higher variable degree generalization of tiled diagonal zipper codes [3] which reduces memory and provides flexibility in hardware implementations.
The proposed codes, like their predecessors in [2, 3], combine aspects of convolutional codes, low-density parity-check (LDPC) codes, and algebraic codes. They can be described as an instance of zipper codes[3] or spatially-coupled generalized LDPC (SC-GLDPC) codes [12]. As a result, they admit a multitude of equivalent descriptions the most natural of which will depend strongly on the background of the reader. Without discounting the potential significance of these connections, we will limit our goal in this paper to providing as simple and self-contained a description of the proposed codes as possible.
The remainder of this paper is organized as follows. We begin in Section II with a special case of our construction corresponding to $L=1$ along with key tools and definitions. In Section III, we apply the resulting construction to obtaining powerful, low-error-floor Hamming-based product-like codes. Finally, we provide the complete construction corresponding to $L\geq 1$ in Section IV, before concluding in Section V with suggestions for future work.
II Generalized Staircase Codes with Higher Bit Degree ($L=1$, $M\geq 1$)
In this section, we provide a generalization of staircase codes [2] in which each coded bit (or nonbinary symbol) is protected by arbitrarily many component codewords rather than only two. The goal is to do this while preserving the property of classical staircase codes, essential to error floor performance, that distinct component codewords intersect in at most one bit. This construction will arise from the combination of a $(1,M)$-DTS and an $(M+1,S)$-net. A $(1,M)$-DTS is equivalent to a Golomb ruler of order $M+1$, to be defined shortly, with the definition of a general $(L,M)$-DTS deferred to Section IV where it is needed.
II-A Code Structure
As with classical staircase codes [2], a codeword is an infinite sequence $B_{0},B_{1},B_{2},\dots$ of $S\times S$ matrices satisfying certain constraints. The parameter $S$ is referred to as the sidelength. Matrix entries are from the binary field $\mathbb{F}_{2}$ for concreteness but nonbinary alphabets are straightforwardly accommodated. We assume zero-based indexing of matrices with the $(i,j)$th entry of a matrix $B$ denoted by $(B)_{(i,j)}$ and the index set $\{0,1,\dots,S-1\}$ denoted by $[S]$.
We begin by specifying a collection of $M+1$ permutations
$\pi_{k}\colon[S]\times[S]$ $\longrightarrow[S]\times[S]$
$(i,j)$ $\longmapsto\pi_{k}(i,j)$
indexed by $k\in[M+1]$ where $M$ is termed the memory parameter. In other words, for each ${k\in[M+1]}$, we have an invertible function $\pi_{k}$ which maps any given $(i,j)\in[S]\times[S]$ to $\pi_{k}(i,j)\in[S]\times[S]$. We further define
$\Pi_{k}\colon\mathbb{F}_{2}^{S\times S}$ $\longrightarrow\mathbb{F}_{2}^{S\times S}$
$B$ $\longmapsto\Pi_{k}(B)$
to be the invertible function mapping a matrix $B\in\mathbb{F}_{2}^{S\times S}$ to its permuted copy $\Pi_{k}(B)\in\mathbb{F}_{2}^{S\times S}$ according to $\pi_{k}$. In particular, the entries of $\Pi_{k}(B)$ are given by
$(\Pi_{k}(B))_{(i,j)}=(B)_{\pi_{k}(i,j)}$
for all $(i,j)\in[S]\times[S]$.
Next, we specify $M+1$ distinct integers
$d_{0}<d_{1}<\dots<d_{M}$
which we refer to as a ruler of order $M+1$. Finally, we specify a linear, systematic $t$-error-correcting component code $\mathcal{C}\subseteq\mathbb{F}_{2}^{(M+1)S}$ of length $(M+1)S$ and dimension $(M+1)S-r$ thus having $r$ parity bits. We can without loss of generality assume that $d_{0}=0$ and that $\Pi_{0}(B)=B$ or, equivalently, $\pi_{0}(i,j)=(i,j)$.
Definition 1 (Weak generalized staircase code).
Given all three of: a collection of permutations, a ruler, and a component code, each as just described, a weak generalized staircase code is defined by the constraint that the rows of the $S\times(M+1)S$ matrix
$\big{(}\Pi_{M}(B_{n-d_{M}})\big{|}\Pi_{M-1}(B_{n-d_{M-1}})\big{|}\cdots\big{|}\Pi_{1}(B_{n-d_{1}})\big{|}B_{n}\big{)}$ (1)
belong to $\mathcal{C}$ for all integers $n\geq d_{M}$ and the initialization condition
$B_{0}=B_{1}=\cdots=B_{d_{M}-1}=0_{S\times S}$ (2)
where $0_{S\times S}$ is the all-zero matrix.
Practically, the first $d_{M}$ blocks (2) need not be transmitted since they are fixed and encoding of the $n$th block is performed recursively starting at $n=d_{M}$. In particular, the $n$th $S\times(S-r)$ block of information bits $B_{n}^{\mathsf{info}}$ produced by a streaming source is adjoined to $M$ permuted previously encoded blocks to form the $S\times((M+1)S-r)$ matrix
$\big{(}\Pi_{M}(B_{n-d_{M}})\big{|}\Pi_{M-1}(B_{n-d_{M-1}})\big{|}\cdots\big{|}\Pi_{1}(B_{n-d_{1}})\big{|}B_{n}^{\mathsf{info}}\big{)}$ (3)
whose rows are encoded using systematic encoders for $\mathcal{C}$ to produce $r$ parity bits for each of the $S$ rows. These parity bits form an $S\times r$ block $B_{n}^{\mathsf{parity}}$ which is adjoined to the information block to produce $B_{n}$ as
$B_{n}=\big{(}\underbrace{\text{---}\;B_{n}^{\mathsf{info}}\;\text{---}}_{S\times(S-r)}\;\big{|}\;\underbrace{\text{--}\;B_{n}^{\mathsf{parity}}\;\text{--}}_{S\times r}\big{)}$
which satisfies the constraint that the rows of (1) belong to $\mathcal{C}$.
Therefore, each transmitted block $B_{n}$ contains $S-r$ information columns and $r$ parity columns yielding a rate of
$R_{\mathsf{unterminated}}=1-\frac{r}{S}$
which we qualify as “unterminated” so that we can later account for rate losses due to the strategy used for ending a transmission. We further observe that we need $r\leq S$ or, equivalently, that the rate of $\mathcal{C}$ is at least $M/(M+1)$.
It remains to specify precisely what the permutations and ruler should be. Before proceeding, we provide a graph perspective which formalizes this question. This can be skimmed over by a reader who is solely interested in a description of the proposed codes without their mathematical justification.
II-B Graph Perspective and Scattering
For mathematical convenience, especially in the later parts of this paper, we extend staircase codewords to be bi-infinite sequences
$\dots,B_{-2},B_{-1},B_{0},B_{1},B_{2},\dots$
where we define
$\cdots=B_{-3}=B_{-2}=B_{-1}=0_{S\times S}$
so that the constraint on the rows of (1) in a generalized staircase code can be presumed to hold for all integers $n\in\mathbb{Z}$ rather than only $n\geq d_{M}$. Practically, nothing changes since the all-zero initialization blocks are not transmitted and only a finite number of past blocks are ever needed for encoding.
We define an infinite bipartite graph whose disjoint and independent vertex sets are $\mathcal{P}=\mathbb{Z}\times[S]\times[S]$ and =×$\mathcal{L}=\mathbb{Z}\times[S]$. An element $(n,i,j)$ of $\mathcal{P}$ corresponds to the $(i,j)$th variable in the $n$th block $(B_{n})_{(i,j)}$ while an element $(n^{\prime},i^{\prime})$ of $\mathcal{L}$ corresponds to the component codeword constraint on the $i^{\prime}$th row of (1) with $n=n^{\prime}$. In particular, the edge set $\mathcal{I}\subseteq\mathcal{P}\times\mathcal{L}$ is given by
=$\mathcal{I}=\big{\{}((n,i,j),(n^{\prime},i^{\prime}))\in\mathcal{P}\times\mathcal{L}\;\big{|}\;\\ n=n^{\prime}-d_{k},k\in[M+1],(i,j)=\pi_{k}(i^{\prime},j^{\prime}),j^{\prime}\in[S]\big{\}}\text{.}$
In other words, this is the (generalized) Tanner graph [13] of the code with $\mathcal{P}$ being the variable nodes and $\mathcal{L}$ being generalized constraint nodes corresponding to membership in the component code $\mathcal{C}$. The variable nodes partition naturally into blocks where the $n$th block comprises $S^{2}$ variable nodes corresponding to $(n,i,j)\in\mathcal{P}$ for every $i,j\in[S]$. Moreover, the constraint nodes partition naturally into constraint spans where the $n^{\prime}$th constraint span corresponds to the $S$ rows of (1) with $n=n^{\prime}$, i.e., contains $S$ constraint nodes corresponding to $(n^{\prime},i^{\prime})\in\mathcal{L}$ for every $i^{\prime}\in[S]$.
We denote by $\mathcal{N}(\mathbf{p})$ the neighbourhood of a given $\mathbf{p}\in\mathcal{P}$ and similarly by $\mathcal{N}(\mathbf{q})$ the neighbourhood of a given $\mathbf{q}\in\mathcal{L}$. Observe that we have for all $(n^{\prime},i^{\prime})\in\mathcal{L}$
$\mathcal{N}((n^{\prime},i^{\prime}))=\{(n,i,j)\in\mathcal{P}\mid n=n^{\prime}-d_{k},k\in[M+1],\\ (i,j)=\pi_{k}(i^{\prime},j^{\prime}),j^{\prime}\in[S]\}$ (4)
with
|$\absolutevalue{\mathcal{N}((n^{\prime},i^{\prime}))}=(M+1)S\text{,}$
i.e., that each constraint involves $(M+1)S$ variables which are a row of (1). Moreover, for all $(n,i,j)\in\mathcal{P}$, we have
$\mathcal{N}((n,i,j))=\{(n^{\prime},i^{\prime})\in\mathcal{L}\mid n^{\prime}=n+d_{k},k\in[M+1],\\ (i^{\prime},j^{\prime})=\pi_{k}^{-1}(i,j),j^{\prime}\in[S]\}$ (5)
with
|$\absolutevalue{\mathcal{N}((n,i,j))}=(M+1)\text{,}$ (6)
i.e., that each variable is involved in $M+1$ constraints.
While each coded symbol is protected by $M+1$ component codewords as per (6), we have no guarantee that our error correcting capability improves with increasing $M$ unless there is some amount of disjointness between the component codewords. One way to guarantee this is by imposing the property that our Tanner graph is $4$-cycle-free. The same property is referred to as scattering in the zipper code framework [3] and as self-orthogonality in the context of threshold decoding [7].
Definition 2 (Scattering/$4$-cycle-freeness/self-orthogonality).
A code with Tanner graph $(\mathcal{P},\mathcal{L},\mathcal{I})$ is scattering if, for all $\mathbf{q}_{1},\mathbf{q}_{2}\in\mathcal{L}$ with $\mathbf{q}_{1}\neq\mathbf{q}_{2}$, we have
|$\absolutevalue{\mathcal{N}(\mathbf{q}_{1})\cap\mathcal{N}(\mathbf{q}_{2})}\leq 1\text{,}$ (7)
i.e., that pairs of distinct constraints share at most one variable.
It is well-known that $4$-cycle-freeness of the Tanner graph of a code has performance implications the specifics of which depend on the channel or error model under consideration as well as the decoding method. In our context, the relevant consequence is as follows:
Proposition 1.
For any code with $4$-cycle-free Tanner graph $(\mathcal{P},\mathcal{L},\mathcal{I})$ having degree $M+1$ variable nodes $\mathcal{P}$ along with generalized constraint nodes $\mathcal{L}$ corresponding to membership in a $t$-error-correcting component code, the Hamming weight of an uncorrectable error under iterative bounded distance decoding of the constraints is at least $(M+1)t+1$.
Proof.
A minimum weight uncorrectable error corresponds to a finite subset of the variables $\mathcal{V}\subseteq\mathcal{P}$ such that any involved constraint in $\mathcal{L}$ is involved in at least $t+1$ of these variables (else, the weight of the error could be further reduced). Given any such error, we can define a finite simple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ in which two vertices are adjacent if and only if they have a common neighbour in the Tanner graph $(\mathcal{P},\mathcal{L},\mathcal{I})$. A fixed vertex $u\in\mathcal{V}$ then has $M+1$ neighbours in $(\mathcal{P},\mathcal{L},\mathcal{I})$ each of which contributes at least $t$ neighbours to $u$ in $\mathcal{G}$. Moreover, these contributions must be disjoint otherwise $4$-cycle-freeness is violated. The degree of $u$ in $\mathcal{G}$ is then at least $(M+1)t$ implying that |$\absolutevalue{\mathcal{V}}\geq(M+1)t+1$. ∎
Our goal is then to find a ruler and permutations which preserve the scattering property (7) which is indeed possessed by classical staircase codes [2]. We first observe that (7) already holds for distinct constraints within the same constraint span since they correspond to different rows of (1). In particular, for all $(n^{\prime},i^{\prime}_{1}),(n^{\prime},i^{\prime}_{2})\in\mathcal{L}$ with $i^{\prime}_{1}\neq i^{\prime}_{2}$, we have
$\mathcal{N}((n^{\prime},i^{\prime}_{1}))\cap\mathcal{N}((n^{\prime},i^{\prime}_{2}))=\emptyset\text{.}$
Therefore, we can restrict our interest to distinct pairs of constraints $(n^{\prime}_{1},i^{\prime}_{1}),(n^{\prime}_{2},i^{\prime}_{2})\in\mathcal{L}$ corresponding to distinct constraint spans $n^{\prime}_{1}\neq n^{\prime}_{2}$ and $i^{\prime}_{1},i_{2}^{\prime}\in[S]$ equal or unequal. Consider now the problem of constructing a point of intersection between these constraint spans $(n,i,j)\in\mathcal{P}$ where
$(n,i,j)\in\mathcal{N}((n^{\prime}_{1},i^{\prime}_{1}))\cap\mathcal{N}((n^{\prime}_{2},i^{\prime}_{2}))$ (8)
and $n^{\prime}_{2}>n^{\prime}_{1}$ without loss of generality. From (4), we see that we must firstly have that
$n=n^{\prime}_{1}-d_{k_{1}}=n^{\prime}_{2}-d_{k_{2}}$
thus
$n^{\prime}_{2}-n^{\prime}_{1}=d_{k_{2}}-d_{k_{1}}>0$ (9)
for some $k_{1},k_{2}\in[M+1]$ with $k_{2}>k_{1}$. Secondly, given such a solution, we must further have that
$(i,j)=\pi_{k_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi_{k_{2}}(i^{\prime}_{2},j^{\prime}_{2})$ (10)
for some $j^{\prime}_{1},j^{\prime}_{2}\in[S]$.
A natural sufficient condition for guaranteeing scattering is then by a choice of ruler so that (9) has a unique solution $k_{1},k_{2}\in[M+1]$ if any and a choice of permutations so that (10) has a unique solution $j^{\prime}_{1},j^{\prime}_{2}\in[S]$. Such solutions determine and thus guarantee the uniqueness of $(n,i,j)$ satisfying (8). We refer to the property of having unique solutions to (9) as inter-block scattering since it is the property that distinct constraint spans share at most one block. On the other hand, we refer to the property of having unique solutions to (10) as intra-block scattering since it is the scattering property restricted to variables within a block. Intra- and inter-block scattering together imply scattering. We now proceed to specify rulers and permutations leading to scattering.
II-C Intra-Block Scattering Permutations via Finite-Geometric Nets
It is both practically convenient and mathematically sufficient to consider permutations defined by linear-algebraic operations on matrix indices $(i,j)\in[S]\times[S]$. To do so, we must associate the elements of the index set $[S]$ with the elements of a finite commutative ring $\mathcal{R}$ of cardinality $S$. A permutation $\pi\colon[S]\times[S]\longrightarrow[S]\times[S]$ can then be interpreted as a linear function $\pi\colon\mathcal{R}\times\mathcal{R}\longrightarrow\mathcal{R}\times\mathcal{R}$ defined by an invertible $2\times 2$ matrix as
$\pi(i,j)=(i,j)\cdot\begin{pmatrix}a&b\\ c&d\end{pmatrix}=(ai+cj,bi+dj)$ (11)
where $a,b,c,d\in\mathcal{R}$ and $(ad-bc)^{-1}$ exists in $\mathcal{R}$. The inverse function is then immediately obtained via the matrix
$(ad-bc)^{-1}\begin{pmatrix}d&-b\\ -c&a\end{pmatrix}\text{.}$
As will be seen shortly, intra-block scattering families of permutations are most easily constructed when $\mathcal{R}$ has sufficiently many elements with invertible differences. Finite fields are an obvious choice since any pair of distinct elements has an invertible difference since all nonzero elements are invertible. We can then take $\mathcal{R}=\mathbb{F}_{q}$, the finite field of order $q$, with $S=q$ a prime power and associate $[S]$ with $\{0,1,\alpha,\alpha^{2},\dots,\alpha^{S-2}\}$ where $\alpha$ is a primitive element of $\mathbb{F}_{q}$. However, a convenient alternative option is to take $\mathcal{R}=\mathbb{Z}_{S}$, the ring of integers modulo $S$, where $S$ is chosen to have a sufficiently large least prime factor denoted $\mathsf{lpf}(S)$. This is because $\{0,1,\dots,\mathsf{lpf}(S)-1\}$ have invertible differences modulo $S$. In particular, we have the following remark:
Remark 1.
For any positive integer $S$ with least prime factor $\mathsf{lpf}(S)$, the integers {±$\{\pm 1,\pm 2,\dots,\pm(\mathsf{lpf}(S)-1)\}$ are coprime with $S$ and are thus multiplicatively invertible modulo $S$.
We now proceed to define $(M+1,S)$-nets which are are well-studied objects in combinatorics and finite geometry with close connections or equivalences to orthogonal arrays, mutually orthogonal Latin squares, transversal designs, and projective planes as discussed in [6].
Definition 3 ($(M+1,S)$-net [6, Part III: Ch. 3]).
An $(M+1,S)$-net is a set of $S^{2}$ elements referred to as points along with a collection of $M+1$ partitions of these points into subsets of size $S$ called lines such that distinct lines intersect in at most one point.
We next make the key observation that $(M+1,S)$-nets are equivalent to intra-block scattering permutations:
Remark 2.
If the (variable-valued) entries of an $S\times S$ matrix $B$ are the points, and the rows of the $M+1$ permuted copies $\Pi_{k}(B)$ for $k\in[M+1]$ are the $M+1$ partitions of these points into lines, then an $(M+1,S)$-net is equivalent to a collection of permutations of $[S]\times[S]$ with the following property: For any distinct $k_{1},k_{2}\in[M+1]$, any row of $\Pi_{k_{1}}(B)$ has a single element in common with any row of $\Pi_{k_{2}}(B)$. Equivalently, for fixed distinct $k_{1},k_{2}\in[M+1]$ and fixed distinct $i^{\prime}_{1},i^{\prime}_{2}\in[S]$, the equation $\pi_{k_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi_{k_{2}}(i^{\prime}_{2},j^{\prime}_{2})$ has a unique solution for $j^{\prime}_{1},j^{\prime}_{2}\in[S]$.
A large number of constructions and open problems are associated with $(M+1,S)$-nets and can be found in [6]. Fortunately, for our practically-oriented purposes, linear-algebraic permutations over conveniently chosen rings already provide enough flexibility and are straightforward to characterize. In particular, we have the following:
Proposition 2 (Complete characterization of linear-algebraic permutations defining $(M+1,S)$-nets).
A collection of $M+1$ permutations of $\mathcal{R}\times\mathcal{R}$, where $\mathcal{R}$ is a finite commutative ring of cardinality S, defined by a collection of invertible $2\times 2$ matrices as in (11), define an $(M+1,S)$-net if and only if, for any pair of distinct matrices $A,\tilde{A}$ in the collection where
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix},\;\tilde{A}=\begin{pmatrix}\tilde{a}&\tilde{b}\\ \tilde{c}&\tilde{d}\end{pmatrix}\text{,}$
we have that $c\tilde{d}-d\tilde{c}$ is invertible in $\mathcal{R}$.
Proof.
This follows by imposing that the linear system $(i,j)A=(\tilde{i},\tilde{j})\tilde{A}$ has a unique solution for $(j,\tilde{j})$ given fixed row indices $(i,\tilde{i})$ by rearranging and imposing an invertible determinant. ∎
We now provide some concrete examples.
Example 1.
If $\mathcal{R}=\mathbb{Z}_{S}$ and $M\leq\mathsf{lpf}(S)$, then the identity permutation $I_{2\times 2}$ together with
$\begin{pmatrix}0&1\\ 1&z\end{pmatrix}$
for $z\in\{0,1,\dots,M-1\}$ define an $(M+1,S)$-net. This can be easily verified by applying Proposition 2 together with Remark 1. Concretely, we can take $\pi_{0}(i,j)=(i,j)$ and define for $k\in\{1,2,\dots,M\}$
$\pi_{k}(i,j)=(j,i+(k-1)j)\pmod{S}$
where $M\leq\mathsf{lpf}(S)$.
Example 2.
If $\mathcal{R}=\mathbb{Z}_{S}$ and $M\leq\mathsf{lpf}(S)$, then the identity permutation $I_{2\times 2}$ together with the involutions
$\begin{pmatrix}-z&1-z^{2}\\ 1&z\end{pmatrix}$
for $z\in\{0,1,\dots,M-1\}$ define an $(M+1,S)$-net. This is verified by noting that Proposition 2 is indifferent to the first rows of the matrices so long as they are invertible. Concretely, we take $\pi_{0}(i,j)=(i,j)$ and define for $k\in\{1,2,\dots,M\}$
$\pi_{k}(i,j)=(-(k-1)i+j,(1-(k-1)^{2})i+(k-1)j)\pmod{S}$
where $M\leq\mathsf{lpf}(S)$.
Example 3.
If $\mathcal{R}=\mathbb{F}_{S}$ with $S=q$ a prime power and $M\leq S$, the identity permutation $I_{2\times 2}$ together with
$\begin{pmatrix}0&1\\ 1&z\end{pmatrix}$
for $z\in\{0,1,\alpha,\alpha^{2},\dots,\alpha^{M-2}\}$ with $\alpha$ a primitive element of $\mathbb{F}_{S}$ form an $(M+1,S)$-net.
We emphasize that in Examples 1 and 2, we must choose $S$ so that $M\leq\mathsf{lpf}(S)$ otherwise an $(M+1,S)$-net is not formed and the error floor performance suffers. If $S=p$ a prime number, then this condition becomes that $M\leq S$. If $M=S$ is required, then Example 3 provides more flexibility in the admissible values of $S$ which can be a prime power.
II-D $(M+1,1)$-nets
It will be useful later to have a family of trivial nets for the case of $S=1$ for all $M\geq 0$. In particular, we interpret the collection of partitions in Definition 3 as allowing for repeated copies of the same partition, or equivalently, repeated copies of the same corresponding permutation. In the case of $S>1$, the definition is unchanged since repeated copies are automatically excluded since they necessarily lead to lines intersecting in $S>1$ points. In the case of $S=1$, we admit $M+1$ copies of the trivial partition of the singleton set into itself leading to lines which intersect in one point by virtue of only containing that point. Alternatively, we define this case explicitly:
Definition 4 ($(M+1,1)$-net).
An $(M+1,1)$-net comprises $M+1$ copies of the trivial partition of a singleton set into itself. The corresponding permutations of $[1]\times[1]$ are $M+1$ copies of the identity permutation $\pi_{k}(i,j)=(i,j)$ for $k\in[M+1]$ and $(i,j)\in[1]\times[1]$, i.e., $(i,j)=(0,0)$.
II-E Inter-Block Scattering via Golomb Rulers
Inter-block scattering rulers are equivalent to Golomb rulers defined as follows:
Definition 5 (Golomb ruler [6, Part VI: Ch. 19]).
A set of $M+1$ distinct integers $d_{0}<d_{1}<\dots<d_{M}$, referred to as marks, form a Golomb ruler of length $d_{M}-d_{0}$ and order $M+1$ if and only if all positive differences between marks $d_{k_{2}}-d_{k_{1}}$ with $k_{2}>k_{1}$ (or distances between distinct marks) are distinct.
We again assume that $d_{0}=0$ without loss of generality since translating a ruler by a constant does not alter its distances. Evidently, for a Golomb ruler, the set of distances or positive differences is the set of positive integers
$\{d_{k_{2}}-d_{k_{1}}\mid k_{2}>k_{1},k_{1},k_{2}\in[M+1]\}$
with cardinality $\binom{M+1}{2}=\frac{(M+1)M}{2}$ implying that this is a lower bound on its length $d_{M}$ which is its largest distance. In particular, we have
$d_{M}\geq\frac{(M+1)M}{2}$ (12)
for any Golomb ruler of order $M+1$ with equality if and only if its set of distances is precisely $\{1,2,\dots,(M+1)M/2\}$.
Remark 3.
Given a Golomb ruler, we see that (9) has a solution whenever $n^{\prime}_{2}-n^{\prime}_{1}$ is in its set of distances and that the solution is unique due to the absence of repeated distances.
Next, observe that Golomb rulers of every order exist by taking, for example, $d_{k}=2^{k}-1$ for $k\in[M+1]$ in which the distance between consecutive marks is larger than the distances between all previous marks. This results in an exponential length $d_{M}=2^{M}-1$ which is very far from the quadratic lower bound (12). Indeed, this is a problem since we can see from (3) that our encoding memory is proportional to the ruler length $d_{M}$. Therefore, we prefer Golomb rulers which have the smallest possible length $d_{M}$ for a given order $M+1$. Such Golomb rulers are referred to as optimal and are well-studied.
We provide for the reader’s convenience optimal Golomb rulers of order up to $M+1=10$ in Table I. Observe that the first four optimal Golomb rulers are verified to be optimal by seeing that they meet the lower bound (12). Beyond these cases, optimal Golomb rulers are typically found by computer search. Optimal Golomb rulers of order up to $M+1=20$ are compiled in [6, Part VI: Ch. 19]. While it is an open problem to exhibit optimal Golomb rulers of every order, Golomb rulers of length $d_{M}<(M+1)^{2}$ have been found for all $M+1<65000$ in [14]. Moreover, an explicit construction of Golomb rulers with cubic length in arbitrary order $M+1$ can also be found in [14].
TABLE I: Some optimal Golomb rulers from [6, Part VI: Ch. 19]
order length optimal Golomb ruler
1 0 0
2 1 0 1
3 3 0 1 3
4 6 0 1 4 6
5 11 0 1 4 9 11
6 17 0 1 4 10 12 17
7 25 0 1 4 10 18 23 25
8 34 0 1 4 9 15 22 32 34
9 44 0 1 5 12 25 27 35 41 44
10 55 0 1 6 10 23 26 34 41 53 55
II-F Recovery of Classical Staircase Codes $(L=1,M=1)$
Consider taking $M=1$ with the optimal Golomb ruler of order $M+1=2$ and using the family of permutations given in Example 1 or in Example 2. We get that $d_{1}=1$ and that $\pi_{1}(i,j)=(j,i)$ or equivalently that $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose). Moreover, for $S>1$, the requirement $M\leq\mathsf{lpf}(S)$ holds trivially since $\mathsf{lpf}(S)\geq 2\geq M$. We therefore recover the classical staircase codes of [2] exactly. This motivates the following definition for a bona fide generalized staircase code:
Definition 6 (Generalized staircase code).
A generalized staircase code is a weak generalized staircase code as in Definition 1 with the added properties that the permutations form an $(M+1,S)$-net and that the ruler is an order $(M+1)$ Golomb ruler.
III Extended-Hamming-Based Generalized Staircase Codes
Staircase codes and related product-like codes [3] typically use a $t$-error-correcting Bose–Chaudhuri–Hocquenghem (BCH) component code $\mathcal{C}$ to protect each coded bit with $M+1=2$ component codewords. This typically requires that $t\geq 3$ to achieve error floors below $10^{-15}$ [15] which imposes a limit on our ability to improve the energy-efficiency of such schemes since the power consumption of a BCH decoder circuit is roughly proportional to $t^{2}$ [16]. However, Proposition 1 suggests that for our generalized staircase codes which are scattering, we can reduce $t$ without increasing the error floor by compensating with a larger bit degree $M+1$. Accordingly, we consider the use of generalized staircase codes with single-error-correcting, i.e., $t=1$, components. Reasoning heuristically, we expect the resulting schemes to be more energy-efficient than those employing a smaller $M+1$ with a larger $t$ since a given correctable error pattern can be corrected with less energy assuming a $t^{2}$ power estimate for the correction of an equal share $t$ of the errors by each of the $M+1$ component codewords. We further anticipate improved performance–complexity–latency tradeoffs in a wider variety of contexts in light of recent works [17, 18] which accomplish such goals via new product-like codes with bit degree increased from two to three.
III-A LUT-Free Decoding of Systematic Hamming Codes
We consider the use of single-error-correcting, double-error-detecting extended Hamming codes which have length $2^{r-1}$ and dimension $2^{r-1}-r$. Such codes have the $r\times 2^{r-1}$ parity-check matrix with columns given by the binary representations of $12j+$ 1 for every $j\in[2^{r-1}]$:
$\begin{pmatrix}0&0&0&\dots&1\\ 0&0&0&\dots&1\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&1\\ 0&0&1&\dots&1\\ 0&1&0&\dots&1\\ 1&1&1&\dots&1\\ \end{pmatrix}\text{.}$ (13)
Famously, syndrome decoding of such codes is trivial with the first $r-1$ bits of the syndrome interpreted as an integer directly giving the error location $j\in[2^{r-1}]$.
Unfortunately, (13) must in general have its columns permuted in order to have a corresponding systematic generator matrix. This would then require that both decoding and generation of columns of the parity-check matrix is done by lookup table (LUT). Both of these operations must be done repeatedly during the iterative decoding of generalized staircase codes. It is thus useful for efficient software-based simulation, and potentially for hardware implementation, to be able to perform these operations as simply as in the case of (13) unpermuted. We accomplish this by means of algebraically-defined systematizing permutations.
In particular, we interpret the column index set $[2^{r-1}]$ as the ring of integers modulo $2^{r-1}$, i.e., $\mathbb{Z}_{2^{r-1}}$. All odd integers are invertible modulo $2^{r-1}$ so that any odd $a\in\mathbb{Z}_{2^{r-1}}$ and any possibly non-odd $b\in\mathbb{Z}_{2^{r-1}}$ define an algebraic permutation $\tau\colon\mathbb{Z}_{2^{r-1}}\longrightarrow\mathbb{Z}_{2^{r-1}}$ where
$\tau(j)=a\cdot j+b\pmod{2^{r-1}}$
and
$\tau^{-1}(j)=a^{-1}\cdot(j-b)\pmod{2^{r-1}}\text{.}$
In the context of software-based simulation, computing $\tau$ and $\tau^{-1}$ can be faster on a modern processor than performing the equivalent table lookup. By a computer search, values of $a$ and $b$ for each $(r-1)\in\{3,4,\dots,15,16\}$ were found such that the the parity-check matrix (13) permuted according to $\tau$ has a systematic generator matrix. The resulting values are provided in Table II. This means that both decoding as well as generating columns of the parity-check matrix can be done as simply as computing $\tau^{-1}$ and $\tau$. Adaptation to a code shortened in the first $s$ positions is accomplished by replacing $b$ with $b+a\cdot s$.
Finally, we report that, somewhat counter-intuitively, the choice of column permutation for the extended Hamming parity-check matrix has a slight but measurable impact on iterative decoding performance even in the absence of shortening. This is because the ordering of the columns impacts the spatial correlations between errors of odd weight greater than or equal to three and the consequent miscorrections. As a result, the use of the extended Hamming codes defined by Table II leads to slightly improved performance relative to systematizing permutations which preserve runs of naturally ordered columns. The same performance gain is also realized by a random choice of systematizing permutation which suggests that this gain is due to a decorrelation effect.
TABLE II: Systematizing algebraic permutations for extended Hamming codes
$r-1$ $a$ $b$ $a^{-1}$
3 1 1 1
4 3 0 11
5 3 0 11
6 3 3 43
7 5 5 77
8 9 11 57
9 19 19 27
10 27 27 531
11 53 53 541
12 89 89 2025
13 163 170 4875
14 301 308 13989
15 553 553 14873
16 1065 1155 55321
III-B Simulations
We consider a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$ which is obtained by shortening a parent extended Hamming code of length $2^{r-1}$ in the first $2^{r-1}-(M+1)S$ positions where ${r-1=\lceil\log_{2}((M+1)S)\rceil}$. The parent codes are as defined by Table II. Furthermore, the permutations are taken to be those of Example 2 thus requiring that $M\leq\mathsf{lpf}(S)$, and the rulers are taken to be the optimal Golomb rulers from Table I. We then assume transmission across a binary symmetric channel (BSC) and perform sliding window decoding of $W$ consecutive blocks at a time where one iteration comprises decoding all constraints in the window consecutively similarly to [2]. To facilitate simulation of the waterfall regime where error propagation events dominate, we convert to a block code via a termination-like strategy in which a frame is formed from $F$ consecutive blocks. For the last $W$ of these $F$ blocks, the information portion is presumed to be zero and only the parity portion is transmitted with the encoder and decoder state being reset between consecutive frames. This results in a block code rate of
$R=\frac{(S-r)(F-W)}{S(F-W)+Wr}$
which approaches $R_{\mathsf{unterminated}}=1-r/S$ as $F$ tends to infinity.
TABLE III: Simulation parameters.
$S$ 669 409 307 179 47
$M$ 3 3 3 4 4
$r$ 13 12 12 11 9
$F$ 725 926 885 1634 912
$W$ (blocks) 21 21 21 36 48
$W$ (Mbits) 9.399 3.513 1.979 1.153 0.106
$R_{\mathsf{unterminated}}$ 0.98057 0.97066 0.96091 0.93855 0.80851
$R$ 0.98000 0.97000 0.96000 0.93725 0.80000
iterations 3 3 4 4 6
gap (dB) 0.585 0.650 0.750 0.950 1.850
input BER 9.86e-4 1.57e-3 2.09e-3 3.25e-3 1.05e-2
Figure 1: Simulation results for extended-Hamming-based generalized staircase codes with parameters given in Table III.
A highly-optimized software-based simulator for the BSC performance of the resulting generalized staircase codes was developed. The simulator is capable of achieving simulated throughputs of several gigabits per second per core of a modern multi-core consumer processor which allows for direct verification of sub-$10^{-15}$ error floors. This is enabled both by the simplicity of syndrome decoding of Hamming components and the fact that decoder behaviour depends only on the syndromes of the received component codewords. Performing Bernoulli sampling of errors for a BSC with crossover probability $\epsilon$ as geometrically-distributed jumps between error locations leads to roughly $\epsilon S^{2}(M+1)$ primitive operations being needed to update the decoder state in response to the reception of a new block. Estimating crudely and heuristically, one decoding iteration similarly takes roughly $\epsilon WS^{2}(M+1)$ primitive operations.
A few representative design examples are considered with parameters given in Table III and the corresponding simulation results given in Fig. 1 as bit error rate (BER) curves. We parameterize the BSC by the equivalent gap to the hard-decision Shannon limit under binary phase-shift keying (BPSK) transmission on an additive white Gaussian noise (AWGN) channel at the respective rates $R$ of the codes considered, rather than by the crossover probability. The last two rows of Table III provide the gap to the hard-decision Shannon limit along with the corresponding BSC crossover probability or input BER for the last point of each curve, i.e., the point of least BER. Where zero bit errors were measured, a dashed line is used in Fig. 1 to indicate the simulated operating point with the number of error-free bit transmissions also indicated. For all but the last code considered, we simulate the transmission of at least $10^{16}$ bits at the last plotted operating point, providing strong confidence that BERs are below $10^{-15}$. Relative to codes using two triple-error-correcting BCH components per bit in [3, 2, 15], the proposed codes perform roughly $0.2$ to $0.5$ dB worse in terms of the gap to the hard-decision Shannon limit as the code rate ranges from $0.98$ down to $0.8$. In return, we expect significantly reduced power consumption. We further see significant reductions in decoding memory thus latency.
IV Generalized Tiled Diagonal Zipper Codes with Higher Bit Degree ($L\geq 1$, $M\geq 1$)
Tiled diagonal zipper codes [3] are a family of scattering generalizations of staircase codes for which the encoding memory approaches half of that of classical staircase codes while maintaining a variable node degree of $M+1=2$. In this section, we further generalize our scattering generalized staircase codes from Section II to both include this family and to extend it to arbitrary bit or variable degree $M+1$. This generalization has two uses. The first is that the encoding and decoding memory can be reduced while keeping all else essentially constant. The second is that it provides flexibility in hardware design by allowing the sidelength to be tuned at a constant code rate. We refer to the resulting scattering code family as higher-order staircase codes since it subsumes and extends the generalized staircase codes of Definition 6.
IV-A Code Structure
Consider starting with a generalized staircase code of sidelength $S^{\prime}$ and memory parameter $M^{\prime}$. We further have our permutations of $[S^{\prime}]\times[S^{\prime}]$ given by $\pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$ and the corresponding matrix-permuting functions given by $\Pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$. Finally, we have a ruler $d_{0}<d_{1}<\dots<d_{M^{\prime}}$ and a component code $\mathcal{C}$ of length $(M^{\prime}+1)S^{\prime}$ and dimension $(M^{\prime}+1)S^{\prime}-r$. However, we have the following twist: We define our constraint that the rows of
$\big{(}\Pi^{\prime}_{M^{\prime}}(B_{n-d_{M^{\prime}}})\big{|}\cdots\big{|}\Pi^{\prime}_{1}(B_{n-d_{1}})\big{|}\Pi^{\prime}_{0}(B_{n-d_{0}})\big{)}$ (14)
belong to $\mathcal{C}$ as holding for all integers $n\in L\mathbb{Z}$ where $L$ is a new fixed parameter. In other words, we restrict the constraint on the rows of (14) to integers $n$ that are multiples of $L$.
Revisiting the Tanner graph perspective, the constraint nodes become $\mathcal{L}=L\mathbb{Z}\times[S]$ and from (5) we see that a variable node $(n,i,j)\in\mathcal{P}$ has the constraint $(n^{\prime},i^{\prime})\in\mathcal{L}$ as a neighbour if and only if $n^{\prime}=n+d_{k}$ for some $k\in[M^{\prime}+1]$ where $n^{\prime}\in L\mathbb{Z}$. Therefore, we get a variable node degree of
$|\absolutevalue{\mathcal{N}((n,i,j))}=\absolutevalue{\{d_{k}\,|\,k\in[M^{\prime}+1],d_{k}\equiv-n\pmod{*}{L}\}}$ | (15)
which depends on the block index $n$ and the ruler. We seek to recover our original variable degree of $M+1$ for all variable nodes independently of $n$. This motivates the following definition:
Definition 7 ($L$-uniform ruler).
An $L$-uniform ruler of order $M^{\prime}+1$ is a set of integers $d_{0}<d_{1}<\dots<d_{M^{\prime}}$ for which
|$\absolutevalue{\{d_{k}\mid k\in[M^{\prime}+1],d_{k}\equiv\ell\pmod{*}{L}\}}=\frac{M^{\prime}+1}{L}$
for each $\ell\in[L]$. Necessarily, $M^{\prime}+1=L(M+1)$ for some positive integer $M+1$ and we can construct any such ruler from $L$ base rulers of order $M+1$ given by
$d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
for each $\ell\in[L]$ as
$\{d_{k}\mid k\in[M^{\prime}+1]\}$ =$=\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1],\ell\in[L]\}$
$=\bigcup_{\ell=0}^{L-1}\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1]\}$ (16)
$=\bigcup_{k=0}^{M}\{Ld^{(\ell)}_{k}+\ell\mid\ell\in[L]\}\text{.}$ (17)
Observe that (16) corresponds to a partition of the ruler into parts which are congruent modulo $L$ while (17) corresponds to a partition into parts which are not congruent modulo $L$. From (15) and (16), we see that an $L$-uniform ruler of order $L(M+1)$ leads to
|$\absolutevalue{\mathcal{N}((n,i,j))}=M+1$
for all $(n,i,j)\in\mathcal{P}$ as needed.
IV-B Scattering Conditions
Next, we observe that the inter-block scattering condition on (9) becomes that for fixed $n^{\prime}_{1},n^{\prime}_{2}\in L\mathbb{Z}$ with $n^{\prime}_{2}>n^{\prime}_{1}$,
$n^{\prime}_{2}-n^{\prime}_{1}=d_{k^{\prime}_{2}}-d_{k^{\prime}_{1}}$
should have a unique solution $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ if any which necessarily satisfies $k_{2}^{\prime}>k^{\prime}_{1}$ and $d_{k^{\prime}_{1}}\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$ since $n^{\prime}_{2}-n^{\prime}_{1}\equiv 0\pmod{*}{L}$. Using an $L$-uniform ruler as in Definition 7, we see that we must have for some $\ell\in[L]$ that $d_{k^{\prime}_{1}},d_{k^{\prime}_{2}}\in\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1]\}$ thus
$n^{\prime}_{2}-n^{\prime}_{1}=d_{k^{\prime}_{2}}-d_{k^{\prime}_{1}}=L\cdot(d^{(\ell)}_{k_{2}}-d^{(\ell)}_{k_{1}})$ (18)
for some $\ell\in[L]$ and $k_{1},k_{2}\in[M+1]$ with $k_{2}>k_{1}$. Therefore, inter-block scattering is obtained when (18) has a unique solution for $\ell\in[L]$ and $k_{1},k_{2}\in[M+1]$ with $k_{2}>k_{1}$.
Secondly, the intra-block scattering condition on (10) becomes that
$\pi^{\prime}_{k^{\prime}_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi^{\prime}_{k^{\prime}_{2}}(i^{\prime}_{2},j^{\prime}_{2})$ (19)
must have a unique solution for $j^{\prime}_{1},j^{\prime}_{2}\in[S^{\prime}]$. However, the uniqueness of solutions to (19) need only be checked for $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ that solve (18) and thus correspond to $d_{k^{\prime}_{1}}\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$. We can then have $\pi^{\prime}_{k^{\prime}_{1}}=\pi^{\prime}_{k^{\prime}_{2}}$ for all $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ such that $d_{k^{\prime}_{1}}\not\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$. This means that we can repeat every permutation $L$ times so that only $(M^{\prime}+1)/L=M+1$ distinct permutations are needed to get intra-block scattering. In particular, consider a collection of $M+1$ permutations of $[S^{\prime}]\times[S^{\prime}]$ given by $\pi_{k}$ for $k\in[M+1]$ and the corresponding matrix-permuting functions given by $\Pi_{k}$ for $k\in[M+1]$. We can take $\pi^{\prime}_{k^{\prime}}=\pi_{k}$ for every $k\in[M+1]$ and $k^{\prime}\in[M^{\prime}+1]$ such that $d_{k^{\prime}}\in\{Ld^{(\ell)}_{k}+\ell\mid\ell\in[L]\}$. Therefore, given $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$, $k_{1},k_{2}\in[M+1]$, and $\ell\in L$ so that $d_{k^{\prime}_{1}}=Ld^{(\ell)}_{k_{1}}+\ell$ and $d_{k^{\prime}_{2}}=Ld^{(\ell)}_{k_{2}}+\ell$, we have $\pi^{\prime}_{k^{\prime}_{1}}=\pi_{k_{1}}$ and $\pi^{\prime}_{k^{\prime}_{2}}=\pi_{k_{2}}$ so that (19) becomes to
$\pi_{k_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi_{k_{2}}(i^{\prime}_{2},j^{\prime}_{2})\text{.}$ (20)
Intra-block scattering then corresponds to uniqueness of solutions to (20) for $j^{\prime}_{1},j^{\prime}_{2}\in[S]$ where $k_{1},k_{2}\in[M+1]$. This is obtained by taking $\pi_{k}$ for $k\in[M+1]$ to correspond to an $(M+1,S^{\prime})$-net.
IV-C Encoding
Next, observe that all base rulers can be translated so that $d_{0}^{(\ell)}=0$ for all $\ell\in[L]$. This results in
$(d_{0},d_{1},\dots,d_{L-1})=(0,1,\dots,L-1)$
for our $L$-uniform ruler $d_{0}<d_{1}<\dots<d_{M^{\prime}}$. This also leads to
$\pi^{\prime}_{0}=\pi^{\prime}_{1}=\cdots=\pi^{\prime}_{L-1}=\pi_{0}$
where $\pi_{0}$ is taken to be the identity permutation. As a result, the constraint span (14) becomes
$\big{(}\Pi^{\prime}_{M^{\prime}}(B_{n-d_{M^{\prime}}})\big{|}\cdots\big{|}\Pi^{\prime}_{L}(B_{n-d_{L}})\big{|}B_{n-(L-1)}\big{|}\cdots\big{|}B_{n-1}\big{|}B_{n}\big{)}$
where $n\in L\mathbb{Z}$. Encoding is performed $L$ blocks at a time with the $(n/L)$th $S^{\prime}\times(LS^{\prime}-r)$ block of information bits $B_{\frac{n}{L}}^{\mathsf{info}}$ adjoined to $M^{\prime}-L+1$ permuted previous blocks to form
$\big{(}\Pi^{\prime}_{M^{\prime}}(B_{n-d_{M^{\prime}}})\big{|}\cdots\big{|}\Pi^{\prime}_{L+1}(B_{n-d_{L+1}})\big{|}\Pi^{\prime}_{L}(B_{n-d_{L}})\big{|}B_{\frac{n}{L}}^{\mathsf{info}}\big{)}$
whose rows are encoded with systematic encoders for the component code $\mathcal{C}$ which is of dimension $(M^{\prime}+1)S^{\prime}-r$ and length $(M^{\prime}+1)S^{\prime}$. This produces an $S^{\prime}\times r$ by block of parity bits $B_{\frac{n}{L}}^{\mathsf{parity}}$ which is adjoined to the information block to produce $L$ consecutive blocks as
$\big{(}B_{n-(L-1)}\big{|}\cdots\big{|}B_{n-1}\big{|}B_{n}\big{)}=\big{(}\underbrace{\text{---}\;B_{\frac{n}{L}}^{\mathsf{info}}\;\text{---}}_{S^{\prime}\times(LS^{\prime}-r)}\;\big{|}\;\underbrace{\text{--}\;B_{\frac{n}{L}}^{\mathsf{parity}}\;\text{--}}_{S^{\prime}\times r}\big{)}$
[…]
To maintain the same code rate and component code $\mathcal{C}$ independently of $L$, we take
$S^{\prime}=\frac{S}{L}$
so that the component code length is $S^{\prime}(M^{\prime}+1)=S(M+1)$ and the code rate is
$R_{\mathsf{unterminated}}=1-\frac{r}{LS^{\prime}}=1-\frac{r}{S}\text{.}$
IV-D Higher-Order Staircase Codes
It remains to define difference triangle sets. In particular, an inter-block scattering $L$-uniform ruler of order $L(M+1)$ arises from an $(L,M)$-DTS defined as follows:
Definition 8 ($(L,M)$-DTS [6, Part VI: Ch. 19]).
An $(L,M)$-DTS is a set of $L$ Golomb rulers of order $M+1$ whose respective sets of distances are disjoint. Equivalently, it is a set of $L$ rulers of order $M+1$ given by
$d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
for $\ell\in[L]$ such that all positive differences
$d^{(\ell)}_{k_{2}}-d^{(\ell)}_{k_{1}}$
for $k_{1},k_{2}\in[M+1]$, $k_{2}>k_{1}$, and $\ell\in[L]$ are distinct.
Evidently, an $L$-uniform ruler of order $M^{\prime}+1=L(M+1)$ whose base rulers are an $(L,M)$-DTS leads to unique solutions to (18) thus inter-block scattering. The construction is summarized by the following definition, which we regard as the central definition, and main contribution, of this paper.
Definition 9 (Higher-order staircase code).
Given all three of the following:
  • an $(L,M)$-DTS given by
    $0=d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
    for $\ell\in[L]$ with corresponding $L$-uniform ruler of order $L(M+1)$
    $d_{0}<d_{1}<\cdots<d_{L(M+1)-1}$
    given accordingly as
    $\{d_{k}\mid k\in[L(M+1)]\}=\\ \{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1],\ell\in[L]\}\text{;}$
  • an $(M+1,S/L)$-net with corresponding $M+1$ permutations of $[S/L]\times[S/L]$ given by $\pi_{k}$ for $k\in[M+1]$ where $\pi_{0}$ is the identity permutation, and a resulting collection of $L(M+1)$ permutations of $[S/L]\times[S/L]$ given by $\pi^{\prime}_{k^{\prime}}=\pi_{k}$ for every $k\in[M+1]$ and $k^{\prime}\in[L(M+1)]$ such that $d_{k^{\prime}}\in\{Ld^{(\ell)}_{k}+\ell\mid\ell\in[L]\}$; and
  • a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$,
a higher-order staircase code of rate $1-r/S$ is defined by the constraint on the bi-infinite sequence of $(S/L)\times(S/L)$ matrices $\dots,B_{-2},B_{-1},B_{0},B_{1},B_{2},\dots$ that the rows of
$\big{(}\Pi^{\prime}_{L(M+1)-1}(B_{n-d_{L(M+1)-1}})\big{|}\cdots\big{|}\Pi^{\prime}_{1}(B_{n-d_{1}})\big{|}\Pi^{\prime}_{0}(B_{n-d_{0}})\big{)}$
belong to $\mathcal{C}$ for all $n\in L\mathbb{Z}$.
We provide a concrete example:
Example 4.
Consider taking $L=2$ and $M=2$. The $L=2$ Golomb rulers of order $M+1=3$ given by
$(d_{0}^{(0)},d_{1}^{(0)},d_{2}^{(0)})$ $=(0,6,7)$
$(d_{0}^{(1)},d_{1}^{(1)},d_{2}^{(1)})$ $=(0,2,5)$
form a $(2,2)$-DTS since their respective sets of distances $\{1,6,7\}$ and $\{2,3,5\}$ are disjoint. We further have
$2\{0,6,7\}$ =$=\{0,12,14\}$
$2\{0,2,5\}+1$ =$=\{1,5,11\}$
thus a corresponding $2$-uniform ruler given by
$(d_{0},d_{1},d_{2},d_{3},d_{4},d_{6})=(0,1,5,11,12,14)\text{.}$
Finally, we have
$\{d_{0},d_{1}\}$ =$=\{2d_{0}^{(0)},2d_{0}^{(1)}+1\}=\{0,1\}$
$\{d_{2},d_{4}\}$ =$=\{2d_{1}^{(0)},2d_{1}^{(1)}+1\}=\{5,12\}$
$\{d_{3},d_{5}\}$ =$=\{2d_{2}^{(0)},2d_{2}^{(1)}+1\}=\{11,14\}$
thus
$\pi^{\prime}_{0}$ $=\pi^{\prime}_{1}=\pi_{0}$
$\pi^{\prime}_{2}$ $=\pi^{\prime}_{4}=\pi_{1}$
$\pi^{\prime}_{3}$ $=\pi^{\prime}_{5}=\pi_{2}\text{.}$
We consider the $(3,S/2)$-net permutations of Example 1 which are given by $\pi_{0}(i,j)=(i,j)$, $\pi_{1}(i,j)=(j,i)$, and $\pi_{2}(i,j)=(j,i+j)\pmod{S/2}$. This leads to $\Pi_{0}(B)=B$, $\Pi_{1}(B)=B^{\mathsf{T}}$, and we further define $B^{\pi}=\Pi_{2}(B)$. Our higher-order staircase code is then defined by the constraint that the rows of
$\big{(}B_{n-14}^{\pi}\big{|}B_{n-12}^{\mathsf{T}}\big{|}B_{n-11}^{\pi}\big{|}B_{n-5}^{\mathsf{T}}\big{|}B_{n-1}\big{|}B_{n}\big{)}$
belong to some specified component code $\mathcal{C}$ for all $n\in 2\mathbb{Z}$.
Figure 2: Visualization of the code of Example 4; rows belong to $\mathcal{C}$.
A visualization of the code of Example 4 is provided in Fig. 2 in the style of zipper codes [3] which better illustrates the encoder structure.
Finally, we note that for higher-order staircase codes as just defined, the code rate, $4$-cycle-freeness, component code, and variable and constraint degrees are all independent of $L$. We therefore expect the threshold and error floor performance to be roughly independent of $L$. For the remainder of this paper, we will be interested in the effect of increasing $L$ while keeping all other parameters fixed.
IV-E Memory
At first glance, encoding requires that memory of the last $d_{M^{\prime}}$ transmitted blocks be maintained. A naive choice would be to maintain the set of $L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}$ blocks $\mathsf{BUFF}_{\frac{n}{L}}$ defined by
$\mathsf{BUFF}_{\frac{n}{L}}=\bigg{\{}B_{n-\mu}\;\bigg{|}\;\\ \mu\in\left\{L,L+1,\dots,L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}+L-1\right\}\bigg{\}}$
for encoding the $(n/L)$th information block. However, by looking at the ruler decomposition (16) we see that we can alternatively maintain $L$ independent buffers $\mathsf{BUFF}^{(\ell)}_{\frac{n}{L}}$ of sizes $d_{M}^{(\ell)}$ for $\ell\in[L]$ given by
$\mathsf{BUFF}^{(\ell)}_{\frac{n}{L}}=\{B_{n-\mu}\mid\mu=L\nu+\ell,\nu\in\{1,2,\dots,d_{M}^{(\ell)}\}\}$
which leads to a potentially reduced memory requirement since
$\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}\leq L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}\text{.}$ (21)
We then have an encoding memory of
$\left(\frac{S}{L}\right)^{2}\cdot\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ (22)
bits (or nonbinary symbols).
On the other hand, the decoding window must typically be chosen as a number of blocks which is a constant multiple of $d_{L(M+1)-1}+1$, the smallest number of consecutive blocks which contains a complete component codeword. If we order our base rulers in descending order of length so that
$\max_{\ell\in[L]}d_{M}^{(\ell)}=d_{M}^{(0)}>d_{M}^{(1)}>\cdots>d_{M}^{(L-1)}$
thus $d_{L(M+1)-1}=L\cdot d_{M}^{(0)}=L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}$, we get
$\left(\frac{S}{L}\right)^{2}\cdot\left(1+L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}\right)$ (23)
as a measure of the decoding memory.
We remark that while (23) is expected to be a reliable proxy for decoding memory assuming that the same trends illustrated in [3] hold for our more general code family, it is not true that (22) is necessarily a reliable proxy for encoding memory. This is because the entirety of every $(S/L)\times(S/L)$ block need not be stored. Instead, a pipelined approach can be employed in which blocks are compressed down to their respective future additive contributions to parity. An example of this can be found in [19] which provides such a hardware implementation of staircase encoders corresponding to $(L=1,M=1)$. However, even with such approaches, there can still be significant hardware efficiency benefits to increasing $L$ as demonstrated in [20] which compares a hardware encoder for a variant of tiled diagonal zipper codes corresponding to $(L=32,M=1)$ to the hardware encoder of [19].
IV-F Notions of Optimality for $(L,M)$-DTS Constructions
As with Golomb rulers which correspond to $(1,M)$-DTSs, it is easy to construct an $(L,M)$-DTS for arbitrary $L$ and $M$ if no further objectives are placed. A well-studied problem is that of constructing an $(L,M)$-DTS of minimum scope where scope is the largest mark across all rulers when rulers are assumed to be translated so that $d_{0}^{(\ell)}=0$ for all $\ell\in[L]$. Under our notation, the scope of an $(L,M)$-DTS is simply $\max_{\ell\in[L]}d_{M}^{(\ell)}$, the largest ruler length. Observe next that we have a trivial lower bound for any $(L,M)$-DTS that
$\max_{\ell\in[L]}d_{M}^{(\ell)}\geq L\cdot\frac{(M+1)M}{2}$ (24)
with equality if and only if the union of the sets of differences of the $L$ rulers is precisely $\{1,2,\dots,L(M+1)M/2\}$.
An $(L,M)$-DTS with minimum scope, or equivalently, minimum maximum length, minimizes decoding memory as per (23). From (23) and (24), we see that the best decoding memory that we can hope to achieve is
$\left(\frac{S}{L}\right)^{2}\left(1+L^{2}\cdot\frac{(M+1)M}{2}\right)=S^{2}\cdot\left(\frac{1}{L^{2}}+\frac{(M+1)M}{2}\right)$
which suggests some decoding memory reduction by increasing $L$.
On the other hand, we see from (22) and (21) that encoding memory is minimized by choice of $(L,M)$-DTS with minimum average length $\frac{1}{L}\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ rather than minimum maximum length $\max_{\ell\in[L]}d_{M}^{(\ell)}$ which leads to two potentially competing notions of optimality. To the best of the authors’ knowledge, the construction of $(L,M)$-DTSs with minimum average length has not been studied. However, in some special cases, an $(L,M)$-DTS with minimum maximum length also has minimum average length. In particular, we have the following:
Proposition 3.
An $(L,M)$-DTS for which the scope (or maximum length) lower bound (24) holds with equality also has minimum average length if $M\in\{0,1,2\}$.
Proof.
In the case of $M=2$, observe that
$2\sum_{\ell=0}^{L-1}d_{2}^{(\ell)}$ $=\sum_{\ell=0}^{L-1}((d_{2}^{(\ell)}-d_{1}^{(\ell)})+d_{1}^{(\ell)}+d_{2}^{(\ell)})$ (25)
$\geq\frac{3L(3L+1)}{2}\text{.}$ (26)
In particular, (25) is a sum over all three distances of all $L$ rulers which must be distinct positive integers. This leads the lower bound (26) which is achieved precisely when the scope lower bound (24) is achieved. In the case of $M=1$, the sum of the lengths is precisely the sum of the distances which means that a similar argument holds while the case of $M=0$ is trivial. ∎
Note that if $M>2$, there is no simple relationship such as (25) between a sum over all ruler lengths and the sum over the sums of their respective distances. Moreover, even in the case of $M=2$, a minimum scope $(L,2)$-DTS which does not meet the lower bound (24) does not minimize average length:
Example 5.
Consider the $(2,2)$-DTS of Example 4 given by $\{\{0,2,5\},\{0,6,7\}\}$ which is of scope $7$. Observe that achieving the scope lower bound of $6$ is impossible in this case since it would require equality of (25) which is even and (26) which is $21$. This implies that $\{\{0,2,5\},\{0,6,7\}\}$ has minimum scope. However, the alternative $(2,2)$-DTS given by $\{\{0,3,4\},\{0,2,7\}\}$ which also has minimum scope also has a smaller average length and thus provides an encoding memory reduction if substituted into Example 4.
IV-G Recovery of Tiled Diagonal Zipper codes $(L\geq 1,M=1)$
Taking $M=1$, a minimum scope $(L,1)$-DTS is given by the rulers $\{0,1\},\{0,2\},\dots,\{0,L\}$ which achieve the lower bound (24) and also minimize average length as per Proposition 3. Consider now the higher-order staircase code defined by this $(L,1)$-DTS and the $(2,S/L)$-nets from Example 1 or Example 2. Taking $d_{1}^{(\ell)}=L-\ell$ for all $\ell\in[L]$, we get
$\{d_{L},d_{L+1},\dots,d_{2L-1}\}$ =$=\{L(L-\ell)+\ell\mid\ell\in[L]\}$
=$=\{L(\ell+1)+L-1-\ell\mid\ell\in[L]\}$
=$=\{L(\ell+2)-(\ell+1)\mid\ell\in[L]\}$
which yields
$\pi^{\prime}_{L}=\pi^{\prime}_{L+1}=\dots=\pi^{\prime}_{2L-1}=\pi_{1}$
where $\pi_{1}(i,j)=(j,i)$. Our higher-order staircase code is then defined by the constraint that the rows of
$\big{(}B_{n-L^{2}}^{\mathsf{T}}\big{|}\cdots\big{|}B_{n-(3L-2)}^{\mathsf{T}}\big{|}B_{n-(2L-1)}^{\mathsf{T}}\big{|}B_{n-(L-1)}\big{|}\cdots\big{|}B_{n-1}\big{|}B_{n}\big{)}$
belong to some specified component code $\mathcal{C}$ for all $n\in L\mathbb{Z}$ which makes these codes identical to the tiled diagonal zipper codes of [3]. Moreover, by (22), we have an encoding memory of
$\left(\frac{S}{L}\right)^{2}\cdot\frac{L(L+1)}{2}=\frac{1}{2}\left(1+\frac{1}{L}\right)S^{2}$ (27)
which approaches half of the encoding memory of classical staircase codes, corresponding to $(L=1,M=1)$, as $L$ grows.
IV-H An Explicitly Solved Case $(L\geq 1,M=2)$
The case of $M=2$ is the only non-trivial instance of the minimum scope $(L,M)$-DTS problem which has an explicitly constructed solution for arbitrary $L$ [21]. In particular, for all $L\equiv 0\pmod{4}$ and $L\equiv 1\pmod{4}$, an explicitly constructed $(L,2)$-DTS of scope $3L$ is given in [22] and this is the minimum scope since it meets the lower bound (24). Furthermore, for all $L\equiv 2\pmod{4}$ and $L\equiv 3\pmod{4}$, an explicitly constructed $(L,2)$-DTS of scope $13L+$ 1 is given in [23] and it is shown that no smaller scope is possible for such $L$ in [22]. Note, however, that [22, 23] consider a different but equivalent problem to the $(L,2)$-DTS problem so some straightforward translations are needed.
Next, note that when $L\equiv 0\pmod{4}$ or $L\equiv 1\pmod{4}$, and only in these cases, the scope lower bound (24) is met which implies by Proposition 3 that the construction of [22] also minimizes average length. In particular, the minimum average length of $43(3L+1)/$ 4 is also achieved. This yields an encoding memory of
$\left(\frac{S}{L}\right)^{2}\cdot\frac{3L(3L+1)}{4}=\frac{3}{4}\left(3+\frac{1}{L}\right)S^{2}$ (28)
which approaches three-quarters of the encoding memory of generalized staircase codes corresponding to $(L=1,M=2)$ as $L$ grows with $L\equiv 0\pmod{4}$ or $L\equiv 1\pmod{4}$. We emphasize that the construction achieving (28) is explicit and is simply Definition 9 together with the explicit $(L,2)$-DTS of [22]. We only omit the description due to its length.
The encoding memory behaviours (27) and (28) indeed show that our higher-order staircase codes naturally generalize the memory reduction property of the tiled diagonal zipper codes of [3] to higher bit degree.
IV-I The General Case $(L\geq 1,M>2)$
In the case of $M>2$, small scope $(L,M)$-DTSs are typically found by combination of algebraic and computer-search methods [21, 24] and specific values of $L$ and $M$ for which a minimum scope $(L,M)$-DTS is known are documented in [6, Part VI: Ch. 19]. The reader is referred to [6] and the references therein for constructions and construction methods. However, as discussed, we may also be interested in minimizing average length as opposed to scope and this can be beneficial as illustrated by Example 5. We suggest this as a problem for future work.
IV-J Recovery of Self-Orthogonal Convolutional Codes ($S=L$)
Letting $S=L$ and recalling Definition 4 for $(M+1,1)$-nets, we see that the net permutations disappear from the construction. Consider further taking $r=1$, i.e., assuming single-parity-check constraints. Definition 9 still remains interesting in this case and leads to a family of rate $(L-1)/L$ codes. These codes are precisely a recursive rather than feedforward version of the self-orthogonal convolutional codes of [7]. In particular, the codes of [7] have a slightly increased rate of $L/(L+1)$ arising from an $(L,M)$-DTS achieved by omitting previous parities in the recursive encoding process. The reader is referred to [25] for an example of how the conversion to a feedforward variant can be done in the case of classical staircase codes. Such conversions typically lead to large error floors since parity symbols are left unprotected.
V Future Work
Several directions for future work are immediately apparent. While we have provided an essentially three-parameter ($L$, $M$, and $t$) family of codes, we have only evaluated performance for $M\geq 1$ with $L=1$ and $t=1$. Most pressing, thus, is the extension of this evaluation to $t\geq 1$ and $L\geq 1$. However, a prerequisite to this is the construction of difference triangle sets having small average length rather than or in addition to small scope. This would enable an analysis of the fundamental performance–complexity–latency tradeoffs of higher-order staircase codes. We may also investigate techniques for analytical error floor estimation and threshold estimation as done in [3] and [26] respectively. Beyond these immediate directions, we may consider irregular-degree and non-scattering variants as done in [17] and [18] respectively.