### v5

# Higher-Order Staircase Codes

Abstract
We generalize staircase codes and tiled diagonal zipper codes, preserving their key properties while allowing each coded symbol to be protected by arbitrarily many component codewords rather than only two. This generalization which we term “higher-order staircase codes” arises from the marriage of two distinct combinatorial objects: difference triangle sets and finite-geometric nets, which have typically been applied separately to code design. We demonstrate one possible realization of these codes, obtaining powerful, high-rate, low-error-floor, and low-complexity coding schemes based on simple iterative syndrome-domain decoding of coupled Hamming component codes. We anticipate that the proposed codes could improve performance–complexity–latency tradeoffs in high-throughput communications applications, most notably fiber-optic, in which classical staircase codes and zipper codes have been applied. We consider the construction of difference triangle sets having minimum scope and sum-of-lengths, which lead to memory-optimal realizations of higher-order staircase codes. These results also enable memory reductions for early families of convolutional codes constructed from difference triangle sets.
Index Terms:
Staircase codes, product codes, finite-geometric nets, difference triangle sets, iterative algebraic decoding.
I Overview
Staircase codes [2] and related product-like codes [3] have provided a highly-competitive paradigm for energy-efficient, high-rate, high-throughput, and high-reliability, error control coding. Such codes use syndrome-domain iterative algebraic or bounded distance decoding of algebraic component codes to enable high decoder throughputs, while maintaining relatively low internal decoder data-flow. Furthermore, their block-convolutional structures translate to natural pipelining and parallelism, the combination of which is essential for efficient hardware designs. Moreover, their typically combinatorial constructions ensure control over error floors which must usually be below $10^{-15}$ in the relevant applications. Most notably, such codes are used in fiber-optic communications [4, Ch. 7], but are also relevant to flash storage applications [5] which increasingly demand similar code characteristics.
In this paper, we provide a fruitful generalization of staircase codes which arises from the combination of difference triangle sets (DTSs) [6, Part VI: Ch. 19] and finite-geometric nets [6, Part III: Ch. 3]. Both DTSs and nets are well-studied combinatorial objects and both have been separately applied to code design, e.g., in [7, 8] and [9, 10] respectively. In these examples, DTSs lead naturally to convolutional codes while nets lead naturally to block codes. It is unsurprising then that their combination leads to a generalization of staircase codes which themselves combine aspects of block and convolutional coding.
We will proceed to summarize at a high level our key construction. A higher-order staircase code is determined by three intertwined objects: a DTS, a net, and a component code. Each of these objects is characterized by two parameters, the meanings of which will be defined later. The relationships between these parameters and properties of the resulting code are as follows: Given an $(L,M)$-DTS, an $(M+1,S/L)$-net, and a component code of length $(M+1)S$ and dimension $(M+1)S-r$, one obtains a rate $1-r/S$ spatially-coupled code in which every symbol is protected by $M+1$ component codewords and any pair of distinct component codewords share at most one symbol. Alternatively, the Tanner graph of the code is an infinite semiregular $4$-cycle-free bipartite graph with variable node degree $M+1$ and generalized constraint node degree $(M+1)S$. The parameter $L$ controls the balance of memory depth against parallelism while keeping all of the aforementioned code properties completely unchanged. A fourth parameter $C$ controls the parallelism obtained by circularly coupling multiple copies of a higher-order staircase code together. The parameters $M$ and $C$ provide mechanisms for improving error floor and threshold performance respectively without altering the component code.
The proposed code family recovers as particular cases some existing code families and interpolates between them, thus providing a unified generalization:
  • When $L=M=C=1$, the classical staircase codes of Smith et al. [2] are recovered.
  • When $L\geq 1$ and $M=C=1$, the tiled diagonal zipper codes of Sukmadji et al. [11] are recovered.
  • When $L\geq 1$, $M=1$, and $C=2$, the OFEC code family [12] is recovered.
  • When $L\geq 1$, $M=1$, and $C\geq 2$, the multiply-chained tiled diagonal zipper codes of Zhao et al. [13] are recovered.
  • When $S/L=M=C=1$, the continuously interleaved codes of Coe [14] are recovered.
  • When $S/L=C=r=1$, a recursive (rather than feedforward) version of the self-orthogonal convolutional codes of Robinson and Bernstein [7] is recovered.
When $M>1$ and $L=1$, we obtain a generalization of staircase codes with higher variable degree which was included in a conference version of this work [1]. A similar construction is included in the recently-issued US patent [15], which recognizes the role of $(1,M)$-DTSs in such a generalization of staircase codes, but makes no explicit use of finite-geometric nets or general $(L,M)$-DTSs. Importantly, when $L>1$ and $M>1$, our construction becomes a higher-variable-degree generalization of tiled diagonal zipper codes [11] which reduces encoding and decoding memory while also providing flexibility in aspects of hardware implementation.
Constructing memory-optimal higher-order staircase codes reduces to a natural extension of the classically studied minimum scope DTS problem [6, Part VI: Ch. 19] in which we additionally seek to minimize a new objective: sum-of-lengths. We study this problem, exhibiting new DTSs that also have implications for earlier work on classical convolutional codes constructed from DTSs.
The proposed codes, like their predecessors in [2, 3], combine aspects of convolutional codes, low-density parity-check (LDPC) codes, and algebraic codes. They can be described as an instance of zipper codes [3], spatially-coupled generalized LDPC (SC-GLDPC) codes [16], or even simply as (highly complex) convolutional codes. However, the proposed codes are very highly structured relative to what is typically admitted under these frameworks. This means that the simplest possible description will be one which is self-contained and does not explicitly appeal to them. Without discounting the potential significance of these connections, we will limit our goal in this paper to providing such a description.
The key advantages of the code structure considered is that it simultaneously enables a large degree of parallelism due to a block-based construction, high performance due to spatial coupling, and control over error floors due to the combinatorial construction. The recovery of established high-throughput coding techniques [2, 11, 12, 14, 13] as special cases is then unsurprising since these properties are precisely what is demanded by such applications.
The remainder of this paper is organized as follows: We begin in Section II with a special case of our construction corresponding to $L=1$ along with key tools and definitions. In Section III, we illustrate one possible realization of the construction of Section II, discussing various practical aspects and a high-throughput simulation technique. Section III also proposes techniques for efficiently decoding systematic Hamming codes without lookup tables (LUTs). Following this, we provide the general, complete construction of higher-order staircase codes corresponding to $L\geq 1$ in Section IV. We follow this with some simulation-based demonstrations in Section V. In Section VI, we study the problem of constructing DTSs that correspond to memory-optimal instances of higher-order staircase codes. A variety of such DTSs are exhibited in the Appendix. Finally, we describe our DTS search technique in Section VII before concluding in Section VIII with suggestions for future work.
II Generalized Staircase Codes with Arbitrary Bit Degree ($L=1$, $M\geq 1$)
In this section, we provide a generalization of staircase codes [2] in which each coded bit (or nonbinary symbol) is protected by arbitrarily many component codewords rather than only two. The goal is to do this while preserving the property of classical staircase codes, essential to error floor performance, that distinct component codewords intersect in at most one bit. This construction will arise from the combination of a $(1,M)$-DTS and an $(M+1,S)$-net. A $(1,M)$-DTS is equivalent to a Golomb ruler of order $M+1$, to be defined shortly, with the definition of a general $(L,M)$-DTS deferred to Section IV where it is needed. Classical staircase codes [2] will be defined as a special case of generalized staircase codes in Section II-F.
II-A Code Structure
As with classical staircase codes [2], a codeword is an infinite sequence $B_{0},B_{1},B_{2},\dots$ of $S\times S$ matrices called blocks satisfying certain constraints. The parameter $S$ is referred to as the sidelength. Matrix entries are from the binary field $\mathbb{F}_{2}$ for concreteness but nonbinary alphabets are straightforwardly accommodated. We assume zero-based indexing of matrices with the $(i,j)$th entry of a matrix $B$ denoted by $(B)_{(i,j)}$ and the index set $\{0,1,\dots,S-1\}$ denoted by $[S]$.
We begin by specifying a collection of $M+1$ permutations
$\pi_{k}\colon[S]\times[S]$ $\longrightarrow[S]\times[S]$
$(i,j)$ $\longmapsto\pi_{k}(i,j)$
indexed by $k\in[M+1]$ where $M$ is termed the degree parameter. In other words, for each ${k\in[M+1]}$, we have an invertible function $\pi_{k}$ which maps any given $(i,j)\in[S]\times[S]$ to $\pi_{k}(i,j)\in[S]\times[S]$. We further define
$\Pi_{k}\colon\mathbb{F}_{2}^{S\times S}$ $\longrightarrow\mathbb{F}_{2}^{S\times S}$
$B$ $\longmapsto\Pi_{k}(B)$
to be the invertible function mapping a matrix $B\in\mathbb{F}_{2}^{S\times S}$ to its permuted copy $\Pi_{k}(B)\in\mathbb{F}_{2}^{S\times S}$ according to $\pi_{k}$. In particular, the entries of $\Pi_{k}(B)$ are given by
$(\Pi_{k}(B))_{(i,j)}=(B)_{\pi_{k}(i,j)}$
for all $(i,j)\in[S]\times[S]$. We will assume that $\Pi_{0}(B)=B$ or, equivalently, that $\pi_{0}(i,j)=(i,j)$, unless otherwise stated. This is typically without loss of generality.
Next, we define rulers:
Definition 1 (Ruler).
A ruler of order $M+1$ is an ordered set of $M+1$ distinct integers $(d_{0},d_{1},\dots,d_{M})$ referred to as marks. A ruler is normalized if $0=d_{0}<d_{1}<\dots<d_{M}$.
We will assume that all rulers are normalized unless otherwise stated. Normalization is achieved by sorting followed by subtraction of the least element from each mark.
Finally, we specify a linear, systematic $t$-error-correcting component code $\mathcal{C}\subseteq\mathbb{F}_{2}^{(M+1)S}$ of length $(M+1)S$ and dimension $(M+1)S-r$ thus having $r$ parity bits. This allows us to provide the following definition for a weak generalized staircase code. The qualifier weak will be dropped once we impose certain conditions on the ruler and permutations.
Definition 2 (Weak generalized staircase code).
Given all three of: a collection of permutations, a ruler, and a component code, each as just described, a weak generalized staircase code is defined by the constraint that the rows of the $S\times(M+1)S$ matrix
$\big{(}\Pi_{M}(B_{n-d_{M}})\big{|}\Pi_{M-1}(B_{n-d_{M-1}})\big{|}\cdots\big{|}\Pi_{1}(B_{n-d_{1}})\big{|}B_{n}\big{)}$ (1)
belong to $\mathcal{C}$ for all integers $n\geq d_{M}$ and the initialization condition
$B_{0}=B_{1}=\cdots=B_{d_{M}-1}=0_{S\times S}$ (2)
where $0_{S\times S}$ is the all-zero matrix.
Practically, the first $d_{M}$ blocks (2) need not be transmitted since they are fixed and encoding of the $n$th block is performed recursively starting at $n=d_{M}$. In particular, the $n$th $S\times(S-r)$ block of information bits $B_{n}^{\mathsf{info}}$ produced by a streaming source is adjoined to $M$ permuted previously encoded blocks to form the $S\times((M+1)S-r)$ matrix
$\big{(}\Pi_{M}(B_{n-d_{M}})\big{|}\Pi_{M-1}(B_{n-d_{M-1}})\big{|}\cdots\big{|}\Pi_{1}(B_{n-d_{1}})\big{|}B_{n}^{\mathsf{info}}\big{)}$ (3)
whose rows are encoded using systematic encoders for $\mathcal{C}$ to produce $r$ parity bits for each of the $S$ rows. These parity bits form an $S\times r$ block $B_{n}^{\mathsf{parity}}$ which is adjoined to the information block to produce $B_{n}$ as
$B_{n}=\big{(}\underbrace{\text{---}\;B_{n}^{\mathsf{info}}\;\text{---}}_{S\times(S-r)}\;\big{|}\;\underbrace{\text{--}\;B_{n}^{\mathsf{parity}}\;\text{--}}_{S\times r}\big{)}$
which satisfies the constraint that the rows of (1) belong to $\mathcal{C}$. Therefore, each transmitted block $B_{n}$ contains $S-r$ information columns and $r$ parity columns yielding a rate of
$R_{\mathsf{unterminated}}=1-\frac{r}{S}$
which we qualify as “unterminated” so that we can later account for rate losses due to the strategy used for ending a transmission. We further observe that we need $r\leq S$ or, equivalently, that the rate of $\mathcal{C}$ is at least $M/(M+1)$.
Figure 1: Visualization of a (classical) staircase code [2] which corresponds to $M=1$, $d_{1}=1$, and $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose); rows belong to $\mathcal{C}$.
Figure 2: Visualization of a (weak) generalized staircase code which corresponds to $M=2$, $(d_{1},d_{2})=(1,3)$, $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose), and $\Pi_{2}(B)=B^{\pi}$ (some permutation); rows belong to $\mathcal{C}$.
We provide a visualization of the resulting codes in Fig. 1 and Fig. 2 for the cases of $M=1$ and $M=2$ respectively in the style of zipper codes [3]. In these pictures, the rightmost column represents the blocks that are actually transmitted while the columns to its left are delayed and permuted copies that are used to constrain the transmitted blocks.
It remains to specify precisely what the permutations and ruler should be. Before proceeding, we provide a graph perspective which formalizes this question. This can be skimmed on a first reading.
II-B Graph Perspective and Scattering
For mathematical convenience, especially in the later parts of this paper, we extend codewords to be bi-infinite sequences
$\dots,B_{-2},B_{-1},B_{0},B_{1},B_{2},\dots$
where we define
$\cdots=B_{-3}=B_{-2}=B_{-1}=0_{S\times S}$
so that the constraint on the rows of (1) in a weak generalized staircase code can be presumed to hold for all integers $n\in\mathbb{Z}$ rather than only $n\geq d_{M}$. Practically, nothing changes since the all-zero initialization blocks are not transmitted and only a finite number of past blocks are ever needed for encoding.
We define an infinite bipartite graph whose disjoint and independent vertex sets are $\mathcal{P}=\mathbb{Z}\times[S]\times[S]$ and =×$\mathcal{L}=\mathbb{Z}\times[S]$. An element $(n,i,j)$ of $\mathcal{P}$ corresponds to the $(i,j)$th variable in the $n$th block $(B_{n})_{(i,j)}$ while an element $(n^{\prime},i^{\prime})$ of $\mathcal{L}$ corresponds to the component codeword constraint on the $i^{\prime}$th row of (1) with $n=n^{\prime}$. In particular, the edge set $\mathcal{I}\subseteq\mathcal{P}\times\mathcal{L}$ is given by
=$\mathcal{I}=\big{\{}((n,i,j),(n^{\prime},i^{\prime}))\in\mathcal{P}\times\mathcal{L}\;\big{|}\;\\ n=n^{\prime}-d_{k},k\in[M+1],(i,j)=\pi_{k}(i^{\prime},j^{\prime}),j^{\prime}\in[S]\big{\}}\text{.}$
In other words, this is the (generalized) Tanner graph [17] of the code with $\mathcal{P}$ being the variable nodes and $\mathcal{L}$ being generalized constraint nodes corresponding to membership in the component code $\mathcal{C}$. The variable nodes partition naturally into blocks where the $n$th block comprises $S^{2}$ variable nodes corresponding to $(n,i,j)\in\mathcal{P}$ for every $i,j\in[S]$. Moreover, the constraint nodes partition naturally into constraint spans where the $n^{\prime}$th constraint span corresponds to the $S$ rows of (1) with $n=n^{\prime}$, i.e., contains $S$ constraint nodes corresponding to $(n^{\prime},i^{\prime})\in\mathcal{L}$ for every $i^{\prime}\in[S]$.
We denote by $\mathcal{N}(\mathbf{p})$ the neighbourhood of a given $\mathbf{p}\in\mathcal{P}$ and similarly by $\mathcal{N}(\mathbf{q})$ the neighbourhood of a given $\mathbf{q}\in\mathcal{L}$. Observe that we have for all $(n^{\prime},i^{\prime})\in\mathcal{L}$
$\mathcal{N}((n^{\prime},i^{\prime}))=\{(n,i,j)\in\mathcal{P}\mid n=n^{\prime}-d_{k},k\in[M+1],\\ (i,j)=\pi_{k}(i^{\prime},j^{\prime}),j^{\prime}\in[S]\}$ (4)
with
|$\absolutevalue{\mathcal{N}((n^{\prime},i^{\prime}))}=(M+1)S\text{,}$
i.e., that each constraint involves $(M+1)S$ variables which are a row of (1). Moreover, for all $(n,i,j)\in\mathcal{P}$, we have
$\mathcal{N}((n,i,j))=\{(n^{\prime},i^{\prime})\in\mathcal{L}\mid n^{\prime}=n+d_{k},k\in[M+1],\\ (i^{\prime},j^{\prime})=\pi_{k}^{-1}(i,j),j^{\prime}\in[S]\}$ (5)
with
|$\absolutevalue{\mathcal{N}((n,i,j))}=(M+1)\text{,}$ (6)
i.e., that each variable is involved in $M+1$ constraints.
While each coded symbol is protected by $M+1$ component codewords as per (6), we have no guarantee that our error correcting capability improves with increasing $M$ unless there is some amount of disjointness between the component codewords. One way to guarantee this is by imposing the property that our Tanner graph is $4$-cycle-free. The same property is referred to as scattering in the zipper code framework [3] and as self-orthogonality in the context of threshold decoding [7].
Definition 3 (Scattering/$4$-cycle-freeness/self-orthogonality).
A code with Tanner graph $(\mathcal{P},\mathcal{L},\mathcal{I})$ is scattering if, for all $\mathbf{q}_{1},\mathbf{q}_{2}\in\mathcal{L}$ with $\mathbf{q}_{1}\neq\mathbf{q}_{2}$, we have
|$\absolutevalue{\mathcal{N}(\mathbf{q}_{1})\cap\mathcal{N}(\mathbf{q}_{2})}\leq 1\text{,}$ (7)
i.e., that pairs of distinct (generalized) constraints share at most one variable.
It is well-known that $4$-cycle-freeness of the Tanner graph of a code has performance implications the specifics of which depend on the channel or error model under consideration as well as the decoding method. In our context, the relevant consequence is as follows:
Proposition 1.
For any code with $4$-cycle-free Tanner graph $(\mathcal{P},\mathcal{L},\mathcal{I})$ having degree $M+1$ variable nodes $\mathcal{P}$ along with generalized constraint nodes $\mathcal{L}$ corresponding to membership in a $t$-error-correcting component code, the Hamming weight of an uncorrectable error under iterative bounded distance decoding of the constraints is at least $(M+1)t+1$.
Proof.
A minimum weight uncorrectable error corresponds to a finite subset of the variables $\mathcal{V}\subseteq\mathcal{P}$ such that any involved constraint in $\mathcal{L}$ is involved in at least $t+1$ of these variables (else, the weight of the error could be further reduced). Given any such error, we can define a finite simple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ in which two vertices are adjacent if and only if they have a common neighbour in the Tanner graph $(\mathcal{P},\mathcal{L},\mathcal{I})$. A fixed vertex $u\in\mathcal{V}$ then has $M+1$ neighbours in $(\mathcal{P},\mathcal{L},\mathcal{I})$ each of which contributes at least $t$ neighbours to $u$ in $\mathcal{G}$. Moreover, these contributions must be disjoint otherwise $4$-cycle-freeness is violated. The degree of $u$ in $\mathcal{G}$ is then at least $(M+1)t$ implying that |$\absolutevalue{\mathcal{V}}\geq(M+1)t+1$. ∎
Note that this bound is usually—but not always—quite loose and does not address the multiplicity of uncorrectable error patterns. Its purpose is solely to provide a theoretical guarantee for asymptotic control over the error floor rather than to estimate the error floor or prescribe parameters.
Our goal is then to find a ruler and permutations which preserve the scattering property (7) which is indeed possessed by classical staircase codes [2]. We first observe that (7) already holds for distinct constraints within the same constraint span since they correspond to different rows of (1). In particular, for all $(n^{\prime},i^{\prime}_{1}),(n^{\prime},i^{\prime}_{2})\in\mathcal{L}$ with $i^{\prime}_{1}\neq i^{\prime}_{2}$, we have
$\mathcal{N}((n^{\prime},i^{\prime}_{1}))\cap\mathcal{N}((n^{\prime},i^{\prime}_{2}))=\emptyset\text{.}$
Therefore, we can restrict our interest to distinct pairs of constraints $(n^{\prime}_{1},i^{\prime}_{1}),(n^{\prime}_{2},i^{\prime}_{2})\in\mathcal{L}$ corresponding to distinct constraint spans $n^{\prime}_{1}\neq n^{\prime}_{2}$ and $i^{\prime}_{1},i_{2}^{\prime}\in[S]$ equal or unequal. Consider now the problem of constructing a point of intersection between these constraint spans $(n,i,j)\in\mathcal{P}$ where
$(n,i,j)\in\mathcal{N}((n^{\prime}_{1},i^{\prime}_{1}))\cap\mathcal{N}((n^{\prime}_{2},i^{\prime}_{2}))$ (8)
and $n^{\prime}_{2}>n^{\prime}_{1}$ without loss of generality. From (4), we see that we must firstly have that
$n=n^{\prime}_{1}-d_{k_{1}}=n^{\prime}_{2}-d_{k_{2}}$
thus
$n^{\prime}_{2}-n^{\prime}_{1}=d_{k_{2}}-d_{k_{1}}>0$ (9)
for some $k_{1},k_{2}\in[M+1]$ with $k_{2}>k_{1}$. Secondly, given such a solution, we must further have that
$(i,j)=\pi_{k_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi_{k_{2}}(i^{\prime}_{2},j^{\prime}_{2})$ (10)
for some $j^{\prime}_{1},j^{\prime}_{2}\in[S]$.
A natural sufficient condition for guaranteeing scattering is then by a choice of ruler so that (9) has a unique solution $k_{1},k_{2}\in[M+1]$ if any and a choice of permutations so that (10) has a unique solution $j^{\prime}_{1},j^{\prime}_{2}\in[S]$. Such solutions determine and thus guarantee the uniqueness of $(n,i,j)$ satisfying (8). We refer to the property of having unique solutions to (9) as inter-block scattering since it is the property that distinct constraint spans share at most one block. On the other hand, we refer to the property of having unique solutions to (10) as intra-block scattering since it is the scattering property restricted to variables within a block. Intra- and inter-block scattering together imply scattering. We now proceed to specify rulers and permutations leading to scattering.
II-C Intra-Block Scattering Permutations via Finite-Geometric Nets
It is both practically convenient and mathematically sufficient to consider permutations defined by linear-algebraic operations on matrix indices $(i,j)\in[S]\times[S]$. To do so, we must associate the elements of the index set $[S]$ with the elements of a finite commutative ring $\mathcal{R}$ of cardinality $S$. A permutation $\pi\colon[S]\times[S]\longrightarrow[S]\times[S]$ can then be interpreted as a linear function $\pi\colon\mathcal{R}\times\mathcal{R}\longrightarrow\mathcal{R}\times\mathcal{R}$ defined by an invertible $2\times 2$ matrix as
$\pi(i,j)=(i,j)\cdot\begin{pmatrix}a&b\\ c&d\end{pmatrix}=(ai+cj,bi+dj)$ (11)
where $a,b,c,d\in\mathcal{R}$ and $(ad-bc)^{-1}$ exists in $\mathcal{R}$. The inverse function is then immediately obtained via the matrix
$(ad-bc)^{-1}\begin{pmatrix}d&-b\\ -c&a\end{pmatrix}\text{.}$
As will be seen shortly, intra-block scattering families of permutations are most easily constructed when $\mathcal{R}$ has sufficiently many elements with invertible differences. Finite fields are an obvious choice since any pair of distinct elements has an invertible difference. We can then take $\mathcal{R}=\mathbb{F}_{q}$, the finite field of order $q$, with $S=q$ a prime power and associate $[S]$ with $\{0,1,\alpha,\alpha^{2},\dots,\alpha^{S-2}\}$ where $\alpha$ is a primitive element of $\mathbb{F}_{q}$. However, a convenient alternative option is to take $\mathcal{R}=\mathbb{Z}_{S}$, the ring of integers modulo $S$, where $S$ is chosen to have a sufficiently large least prime factor denoted $\mathsf{lpf}(S)$. This is because $\{0,1,\dots,\mathsf{lpf}(S)-1\}$ have invertible differences modulo $S$. In particular, we have the following remark:
Remark 1.
For any positive integer $S$ with least prime factor $\mathsf{lpf}(S)$, the integers {±$\{\pm 1,\pm 2,\dots,\pm(\mathsf{lpf}(S)-1)\}$ are coprime with $S$ and are thus multiplicatively invertible modulo $S$.
We now proceed to define $(M+1,S)$-nets which are well-studied objects in combinatorics and finite geometry with close connections or equivalences to orthogonal arrays, mutually orthogonal Latin squares, transversal designs, and projective planes as discussed in [6].
Definition 4 ($(M+1,S)$-net [6, Part III: Ch. 3]).
An $(M+1,S)$-net is a set of $S^{2}$ elements referred to as points along with a collection of $M+1$ partitions of these points into subsets of size $S$ called lines such that distinct lines intersect in at most one point.
We next make the key observation that $(M+1,S)$-nets are equivalent to intra-block scattering permutations:
Remark 2.
If the (variable-valued) entries of an $S\times S$ matrix $B$ are the points, and the rows of the $M+1$ permuted copies $\Pi_{k}(B)$ for $k\in[M+1]$ are the $M+1$ partitions of these points into lines, then an $(M+1,S)$-net is equivalent to a collection of permutations of $[S]\times[S]$ with the following property: For any distinct $k_{1},k_{2}\in[M+1]$, any row of $\Pi_{k_{1}}(B)$ has a single element in common with any row of $\Pi_{k_{2}}(B)$. Equivalently, for fixed distinct $k_{1},k_{2}\in[M+1]$ and fixed distinct $i^{\prime}_{1},i^{\prime}_{2}\in[S]$, the equation $\pi_{k_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi_{k_{2}}(i^{\prime}_{2},j^{\prime}_{2})$ has a unique solution for $j^{\prime}_{1},j^{\prime}_{2}\in[S]$.
A large number of constructions and open problems are associated with $(M+1,S)$-nets and can be found in [6]. Fortunately, for our practically-oriented purposes, linear-algebraic permutations over conveniently chosen rings already provide enough flexibility and are trivial to characterize. In particular, we have the following:
Proposition 2 (Complete characterization of linear-algebraic permutations defining $(M+1,S)$-nets).
A collection of $M+1$ permutations of $\mathcal{R}\times\mathcal{R}$, where $\mathcal{R}$ is a finite commutative ring of cardinality S, defined by a collection of invertible $2\times 2$ matrices as in (11), define an $(M+1,S)$-net if and only if, for any pair of distinct matrices $A,\tilde{A}$ in the collection where
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix},\;\tilde{A}=\begin{pmatrix}\tilde{a}&\tilde{b}\\ \tilde{c}&\tilde{d}\end{pmatrix}\text{,}$
we have that $c\tilde{d}-d\tilde{c}$ is invertible in $\mathcal{R}$.
Proof.
This follows by imposing that the linear system $(i,j)A=(\tilde{i},\tilde{j})\tilde{A}$ has a unique solution for $(j,\tilde{j})$ given fixed row indices $(i,\tilde{i})$ by rearranging and imposing an invertible determinant. ∎
We now provide some concrete examples.
Example 1.
If $\mathcal{R}=\mathbb{Z}_{S}$ and $M\leq\mathsf{lpf}(S)$, then the identity permutation $I_{2\times 2}$ together with
$\begin{pmatrix}0&1\\ 1&z\end{pmatrix}$
for $z\in\{0,1,\dots,M-1\}$ define an $(M+1,S)$-net. This can be easily verified by applying Proposition 2 together with Remark 1. Concretely, we can take $\pi_{0}(i,j)=(i,j)$ and define for $k\in\{1,2,\dots,M\}$
$\pi_{k}(i,j)=(j,i+(k-1)j)\pmod{S}$
where $M\leq\mathsf{lpf}(S)$.
Example 2.
If $\mathcal{R}=\mathbb{Z}_{S}$ and $M\leq\mathsf{lpf}(S)$, then the identity permutation $I_{2\times 2}$ together with the involutions
$\begin{pmatrix}-z&1-z^{2}\\ 1&z\end{pmatrix}$
for $z\in\{0,1,\dots,M-1\}$ define an $(M+1,S)$-net. This is verified by noting that Proposition 2 is indifferent to the first rows of the matrices so long as they are invertible. Concretely, we take $\pi_{0}(i,j)=(i,j)$ and define for $k\in\{1,2,\dots,M\}$
$\pi_{k}(i,j)=(-(k-1)i+j,(1-(k-1)^{2})i+(k-1)j)\pmod{S}$
where $M\leq\mathsf{lpf}(S)$.
Example 3.
If $\mathcal{R}=\mathbb{F}_{S}$ with $S=q$ a prime power and $M\leq S$, the identity permutation $I_{2\times 2}$ together with
$\begin{pmatrix}0&1\\ 1&z\end{pmatrix}$
for $z\in\{0,1,\alpha,\alpha^{2},\dots,\alpha^{M-2}\}$ with $\alpha$ a primitive element of $\mathbb{F}_{S}$ form an $(M+1,S)$-net.
We emphasize that in Examples 1 and 2, we must choose $S$ so that $M\leq\mathsf{lpf}(S)$ otherwise an $(M+1,S)$-net is not formed. If $S=p$ a prime number, then this condition becomes that $M\leq S$. If $M=S$ is required, then Example 3 provides more flexibility in the admissible values of $S$ which can be a prime power.
II-D $(M+1,1)$-nets
It will be useful later to have a family of trivial nets for the case of $S=1$ for all $M\geq 0$. In particular, we interpret the collection of partitions in Definition 4 as allowing for repeated copies of the same partition, or equivalently, repeated copies of the same corresponding permutation. In the case of $S>1$, the definition is unchanged since repeated copies are automatically excluded since they necessarily lead to lines intersecting in $S>1$ points. In the case of $S=1$, we admit $M+1$ copies of the trivial partition of the singleton set into itself leading to lines which intersect in one point by virtue of only containing that point. Alternatively, we define this case explicitly:
Definition 5 ($(M+1,1)$-net).
An $(M+1,1)$-net comprises $M+1$ copies of the trivial partition of a singleton set into itself. The corresponding permutations of $[1]\times[1]$ are $M+1$ copies of the identity permutation $\pi_{k}(i,j)=(i,j)$ for $k\in[M+1]$ and $(i,j)\in[1]\times[1]$, i.e., $(i,j)=(0,0)$.
II-E Inter-Block Scattering via Golomb Rulers
Inter-block scattering rulers are equivalent to Golomb rulers defined as follows:
Definition 6 (Golomb ruler [6, Part VI: Ch. 19]).
An order $M+1$ ruler $(d_{0},d_{1},\dots,d_{M})$ is a Golomb ruler if all differences between distinct marks $d_{k_{1}}-d_{k_{2}}$ where $k_{1},k_{2}\in[M+1]$ and $k_{1}\neq k_{2}$, are distinct. The length of a Golomb ruler is its largest such difference.
Unless otherwise stated, we will assume that all Golomb rulers are normalized so that $0=d_{0}<d_{1}<\dots<d_{M}$. The length of a Golomb ruler is then simply its largest mark $d_{M}$. The set of distances of a Golomb ruler is the set of its positive differences
$\{d_{k_{2}}-d_{k_{1}}\mid k_{2}>k_{1},k_{1},k_{2}\in[M+1]\}$
which is necessarily of cardinality $\binom{M+1}{2}=\frac{(M+1)M}{2}$ by the distinctness of these differences. This implies that the length $d_{M}$ of an order $M+1$ Golomb ruler satisfies the trivial bound
$d_{M}\geq\frac{(M+1)M}{2}$ (12)
with equality if and only if its set of distances is precisely $\{1,2,\dots,(M+1)M/2\}$. A Golomb ruler for which the trivial bound holds with equality is said to be perfect.
Remark 3.
Given a Golomb ruler, we see that (9) has a solution whenever $n^{\prime}_{2}-n^{\prime}_{1}$ is in its set of distances and that the solution is unique due to the absence of repeated distances.
Next, observe that Golomb rulers of every order exist by taking, for example, $d_{k}=2^{k}-1$ for $k\in[M+1]$ in which the distance between consecutive marks is larger than the distances between all previous marks. This results in an exponential length $d_{M}=2^{M}-1$ which is very far from the quadratic lower bound (12). Indeed, this is a problem since we can see from (3) that our encoding memory is proportional to the ruler length $d_{M}$. Therefore, we prefer Golomb rulers which have the smallest possible length $d_{M}$ for a given order $M+1$. Such Golomb rulers are referred to as optimal and are well-studied.
We provide for the reader’s convenience optimal Golomb rulers of order up to $M+1=10$ in Table I. Observe that the first four optimal Golomb rulers are verified to be optimal by seeing that they are perfect. Beyond these cases, optimal Golomb rulers are typically found by computer search. Optimal Golomb rulers of order up to $M+1=20$ are compiled in [6, Part VI: Ch. 19]. While it is an open problem to exhibit optimal Golomb rulers of every order, Golomb rulers of length $d_{M}<(M+1)^{2}$ have been found for all $M+1<65000$ in [18]. Moreover, an explicit construction of Golomb rulers with cubic length in arbitrary order $M+1$ can also be found in [18].
TABLE I: Some optimal Golomb rulers from [6, Part VI: Ch. 19]
order length optimal Golomb ruler
1 0 0
2 1 0 1
3 3 0 1 3
4 6 0 1 4 6
5 11 0 1 4 9 11
6 17 0 1 4 10 12 17
7 25 0 1 4 10 18 23 25
8 34 0 1 4 9 15 22 32 34
9 44 0 1 5 12 25 27 35 41 44
10 55 0 1 6 10 23 26 34 41 53 55
II-F Recovery of Classical Staircase Codes $(L=1,M=1)$
Consider taking $M=1$ with the optimal Golomb ruler of order $M+1=2$ and using the family of permutations given in Example 1 or in Example 2. We get that $d_{1}=1$ and that $\pi_{1}(i,j)=(j,i)$ or, equivalently, that $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose). Moreover, for $S>1$, the requirement $M\leq\mathsf{lpf}(S)$ holds trivially since $\mathsf{lpf}(S)\geq 2\geq M$. We therefore recover the classical staircase codes of [2] exactly. This motivates the following definition for a bona fide generalized staircase code:
Definition 7 (Generalized staircase code).
A generalized staircase code is a weak generalized staircase code as in Definition 2 with the added properties that the permutations form an $(M+1,S)$-net and that the ruler is an order $M+1$ Golomb ruler.
III Extended-Hamming-Based Generalized Staircase Codes
Staircase codes and related product-like codes [3] typically use a $t$-error-correcting Bose–Ray-Chaudhuri–Hocquenghem (BCH) component code $\mathcal{C}$ to protect each coded bit with $M+1=2$ component codewords. This typically requires that $t\geq 3$ to achieve error floors below $10^{-15}$ [19] which imposes a limit on our ability to improve the energy-efficiency of such schemes since the energy cost of a BCH decoding is roughly $\mathcal{O}(t^{2})$ [20]. However, Proposition 1 suggests that for our generalized staircase codes which are scattering, we can reduce $t$ without increasing the error floor by compensating with a larger bit degree $M+1$. Accordingly, we consider the use of generalized staircase codes with single-error-correcting, i.e., $t=1$, components. Reasoning heuristically, we expect the resulting schemes to be more energy-efficient than those employing a smaller $M+1$ with a larger $t$ since the energy cost per corrected bit would be $\mathcal{O}(t^{2}/t)=\mathcal{O}(t)$ with the lower bound of Proposition 1 held constant.
We consider the use of single-error-correcting, double-error-detecting extended Hamming codes which have length $2^{r-1}$ and dimension $2^{r-1}-r$. Such codes have the $r\times 2^{r-1}$ parity-check matrix with columns given by the binary representations of $12j+$ 1 for every $j\in[2^{r-1}]$:
$\begin{pmatrix}0&0&0&0&\cdots&1\\ 0&0&0&0&\cdots&1\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&1\\ 0&0&1&1&\cdots&1\\ 0&1&0&1&\cdots&1\\ 1&1&1&1&\cdots&1\end{pmatrix}\text{.}$ (13)
Famously, syndrome decoding of such codes is trivial with the first $r-1$ bits of the syndrome interpreted as an integer directly giving the error location $j\in[2^{r-1}]$.
Unfortunately, (13) must in general have its columns permuted in order to have a corresponding systematic generator matrix. This would then require that both decoding and generation of columns of the parity-check matrix is done by lookup table (LUT). Both of these operations must be done repeatedly during the iterative decoding of generalized staircase codes. It is thus useful, both for efficient software-based simulation and for hardware implementation, to be able to perform these operations as simply as in the case of (13) unpermuted. We provide some solutions to this problem for software and hardware respectively in Sections III-B and III-C.
III-A Simulations
We consider a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$ which is obtained by shortening a parent extended Hamming code of length $2^{r-1}$ in the first $2^{r-1}-(M+1)S$ positions where ${r-1=\lceil\log_{2}((M+1)S)\rceil}$. The parent codes are as defined by Table V later. Furthermore, the permutations are taken to be those of Example 2 thus requiring that $M\leq\mathsf{lpf}(S)$, and the rulers are taken to be the optimal Golomb rulers from Table I. We then assume transmission across a binary symmetric channel (BSC) and perform sliding window decoding of $W$ consecutive blocks at a time where one iteration comprises decoding all constraints in the window consecutively and $I$ iterations are performed before advancing the window by one block similarly to [2]. To facilitate simulation of the waterfall regime where error propagation events dominate, we convert to a block code via a termination-like strategy in which a frame is formed from $F$ consecutive blocks. For the last $W$ of these $F$ blocks, the information portion is presumed to be zero and only the parity portion is transmitted with the encoder and decoder state being reset between consecutive frames. This results in a block code rate of
$R=\frac{(S-r)(F-W)}{S(F-W)+Wr}$
which approaches $R_{\mathsf{unterminated}}=1-r/S$ as $F\to\infty$. The reader is referred to [3] for alternative decoding and termination-like procedures for similar families of codes.
Figure 3: Simulation results for extended-Hamming-based generalized staircase codes with parameters given in Table II.
TABLE II: Simulation parameters for Fig. 3.
$S$ $M$ $F$ $W$ (Mbits) $R$ $I$ gap (dB) in. BER
669 3 725 21 (9.40) 0.980 3 0.585 9.86e-4
409 3 926 21 (3.51) 0.970 3 0.650 1.57e-3
307 3 885 21 (1.98) 0.960 4 0.750 2.09e-3
307 3 717 17 (1.60) 0.960 4 0.750 2.09e-3
179 4 1634 36 (1.15) 0.937 4 0.950 3.25e-3
179 4 1089 24 (0.77) 0.937 4 0.950 3.25e-3
47 4 912 48 (0.11) 0.800 6 1.850 1.05e-2
We developed a highly-optimized software simulator for the BSC performance of the resulting extended-Hamming-based generalized staircase codes. This simulator is available online [21] and is capable of achieving simulated throughputs of several gigabits per second per core of a modern multi-core consumer processor which allows for direct verification of sub-$10^{-15}$ error floors. This is enabled both by the simplicity of syndrome decoding of Hamming components and the fact that decoder behaviour depends only on the syndromes of the received component codewords. Therefore, simulation can be performed directly in syndrome domain with individual syndromes represented by individual integers. Performing Bernoulli sampling of errors for a BSC with crossover probability $\epsilon$ as geometrically-distributed jumps between error locations leads to roughly $\epsilon S^{2}(M+1)$ primitive operations being needed to update the decoder state in response to the reception of a new block. Estimating crudely and heuristically, one decoding iteration similarly takes roughly $\epsilon WS^{2}(M+1)$ primitive operations.
A few representative design examples are considered with parameters given in Table II and the corresponding simulation results given in Fig. 3 as bit error rate (BER) curves. We parameterize the BSC by the equivalent gap to the hard-decision Shannon limit under binary phase-shift keying (BPSK) transmission on an additive white Gaussian noise (AWGN) channel at the respective rates $R$ of the codes considered, rather than by the crossover probability. The last two rows of Table II provide the gap to the hard-decision Shannon limit along with the corresponding BSC crossover probability or input BER for the last point of each curve, i.e., the point of least BER. Where zero bit errors were measured, a dashed line is used in Fig. 3 to indicate the simulated operating point with the number of error-free bit transmissions also indicated. For all but the last code considered, we simulate the transmission of at least $10^{16}$ bits at the last plotted operating point, providing strong confidence that BERs are below $10^{-15}$.
Relative to codes using two triple-error-correcting BCH components per bit in [3, 2, 19], the example codes perform roughly $0.2$ to $0.5$ dB worse in terms of the gap to the hard-decision Shannon limit as the code rate ranges from $0.98$ down to $0.8$. In return, we get reduced complexity and latency. Some heuristic estimates of performance, latency, and complexity in comparison to classical staircase codes [2] are provided in Tables III and IV. In high-throughput contexts, latency is directly proportional to the number of bits in the decoding window $WS^{2}$. On the other hand, the number of component codewords in the decoding window is about $WS$. In practice, only a fraction of these must be decoded each iteration and we assume this fraction to be constant for simplicity. Moreover, we assume that for the relatively small values of $M$ considered, the cost of each component code decoding is dominated by the cost of invoking the component code decoder $t^{2}$ rather than the cost of updating the syndromes of codewords involved in each corrected bit. This leads to a heuristic complexity measure of $IWSt^{2}$.
TABLE III: Heuristic estimates of performance, latency, and complexity for generalized staircase codes with $R\approx 0.937$
staircase [2, 22] this work
gap to Shannon at 1e-15 0.56 dB 0.95 dB
latency: $WS^{2}$ 1.8e6 7.7e5
iterations: $I$ 3 4
decodings per iteration: $WS$ 3.6e3 4.3e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWSt^{2}$ 9.7e4 1.7e4
TABLE IV: Heuristic estimates of performance, latency, and complexity for generalized staircase codes with $R\approx 0.96$
staircase [2, 23] this work
gap to Shannon at 1e-15 0.49 dB 0.75 dB
latency: $WS^{2}$ 4e6 1.6e6
iterations: $I$ 4 4
decodings per iteration: $WS$ 5e3 5.2e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWSt^{2}$ 18e4 2e4
III-B LUT-Free Decoding of Systematic Hamming Codes with Affine Permutations in $\mathbb{Z}_{2^{r-1}}$
We now revisit the problem of constructing systematic Hamming codes that admit a LUT-free decoding. This amounts to finding a permutation of the columns of (13) such that a corresponding systematic generator matrix exists and both the forward and inverse permutations are easily calculated without a LUT. Note that the existence of a systematic generator matrix is equivalent to the invertibility of the $r\times r$ matrix formed by the last $r$ columns of the parity-check matrix. We will refer to a permutation of the columns of (13) achieving this as a systematizing permutation.
A natural approach to this problem is to consider affine permutations in $\mathbb{Z}_{2^{r-1}}$. In particular, we interpret the column index set $[2^{r-1}]$ as the ring of integers modulo $2^{r-1}$, i.e., $\mathbb{Z}_{2^{r-1}}$. All odd integers are invertible modulo $2^{r-1}$ so that any odd $a\in\mathbb{Z}_{2^{r-1}}$ and any possibly non-odd $b\in\mathbb{Z}_{2^{r-1}}$ define an algebraic permutation $\tau\colon\mathbb{Z}_{2^{r-1}}\longrightarrow\mathbb{Z}_{2^{r-1}}$ where
$\tau(j)=aj+b\pmod{2^{r-1}}$
and
$\tau^{-1}(j)=a^{-1}(j-b)\pmod{2^{r-1}}\text{.}$
In the context of software-based simulation, computing $\tau$ and $\tau^{-1}$ can be faster on a modern processor than performing the equivalent table lookup. By a computer search, values of $a$ and $b$ for each $(r-1)\in\{3,4,\dots,16\}$ were found such that $\tau$ is systematizing. The resulting values are provided in Table V. This means that both decoding a syndrome as well as generating columns of the parity-check matrix can be done as simply as computing $\tau^{-1}$ and $\tau$. Adaptation to a code shortened in the first $s$ positions is accomplished by replacing $b$ with $b+as$.
We report that, somewhat counter-intuitively, the choice of column permutation for the extended Hamming parity-check matrix leads to a slight but measurable improvement in iterative decoding performance even in the absence of shortening. This is because the ordering of the columns impacts the spatial correlations between errors of odd weight greater than or equal to three and the consequent miscorrections. As a result, the use of the extended Hamming codes defined by Table V leads to slightly improved performance relative to systematizing permutations which preserve runs of naturally ordered columns. The same performance gain is also realized by a random choice of systematizing permutation which suggests that this gain is due to a decorrelation effect.
TABLE V: Parameters for systematizing affine permutations for extended Hamming codes
$r-1$ $a$ $b$ $a^{-1}$
3 1 1 1
4 3 0 11
5 3 0 11
6 3 3 43
7 5 5 77
8 9 11 57
9 19 19 27
10 27 27 531
11 53 53 541
12 89 89 2025
13 163 170 4875
14 301 308 13989
15 553 553 14873
16 1065 1155 55321
III-C LUT-Free Decoding of Systematic Hamming Codes with Boolean Functions
TABLE VI: Parameters for systematizing permutations for extended Hamming Codes via Boolean Functions
$i$ $d^{(i)}$ $\sum_{u=0}^{d^{(i)}-1}s^{(i)}_{u}2^{u}$
2 2 0
3 2 0
4 3 0
5 3 1
6 3 2
7 2 0
8 4 0
9 4 1
10 4 2
11 4 3
12 4 4
13 4 5
14 4 6
15 3 0
In the context of software implementation, the approach presented in Section III-B is as good as we can hope for since modern processors already have dedicated multipliers. On the other hand, in the context of hardware implementation, we might wish to avoid realizing multipliers—which are expensive—altogether and directly seek a permutation calculable with a small number of logic gates. To this end, we extend an approach to LUT-free Hamming decoding provided in [24].
Firstly, we interpret the column index set $[2^{r-1}]$ as the vector space $\mathbb{F}_{2}^{r-1}$ taking $j\in[2^{r-1}]$ with binary representation $j=\sum_{u=0}^{r-2}b_{u}2^{u}$ with $b_{u}\in\{0,1\}$ to correspond to the binary vector $(b_{0},b_{1},\ldots,b_{r-2})\in\mathbb{F}_{2}^{r-1}$.
Consider the class of permutations of $\mathbb{F}_{2}^{r-1}$ which map $(b_{0},b_{1},\dots,b_{r-2})\in\mathbb{F}_{2}^{r-1}$ to $(c_{0},c_{1},\dots,c_{r-2})\in\mathbb{F}_{2}^{r-1}$ and take the form
$b_{0}$ $\mapsto c_{0}=b_{0}$
$b_{1}$ $\mapsto c_{1}=b_{1}$
$b_{i}$ $\mapsto c_{i}=b_{i}+\prod_{u=0}^{d^{(i)}-1}(s^{(i)}_{u}+b^{(i)}_{u})\;\text{for $2\leq i\leq r-2$}$
where $d^{(i)}$ and $(s^{(i)}_{0},s^{(i)}_{1},\dots,s^{(i)}_{d^{(i)}-1})$ for $22\leq i\leq r-$ 2 are parameters defining the permutation. We further impose that $d^{(i)}\leq i$ which guarantees invertibility since the expression
$c_{i}=b_{i}+\prod_{u=0}^{d^{(i)}-1}(s^{(i)}_{u}+b^{(i)}_{u})$
allows the recovery of $b_{i}$ from $c_{i}$ and $b_{0},b_{1},\dots,b_{i-1}$ and thus all $b_{i}$ from $c_{i}$ by induction. We will represent $(s^{(i)}_{0},s^{(i)}_{1},\dots,s^{(i)}_{d^{(i)}-1})$ with the corresponding integer $\sum_{u=0}^{d^{(i)}-1}s^{(i)}_{u}2^{u}$ for convenience.
This class of permutations generalizes the specific example given in [24] for $r=8$ and is easy to implement in hardware requiring for each $i$ a $d^{(i)}$-input AND gate, one XOR gate, and a NOT gate for every nonzero value of $s^{(i)}_{u}$.
Table VI defines a special choice of systematizing permutations for extended Hamming codes of length up to $2^{16}$ with the property that the table can be truncated at any row while maintaining the systematizing property. These were found by a computer search. For example, for extended Hamming codes of length $642^{6}=$ 64, the map defined by Table VI is
$(b_{0},b_{1},b_{2},b_{3},b_{4},b_{5})\mapsto\\ (b_{0},b_{1},b_{2}+b_{0}b_{1},b_{3}+b_{0}b_{1},b_{4}+b_{0}b_{1}b_{2},b_{5}+(1+b_{0})b_{1}b_{2})$
The last seven columns of our parity-check matrix are are the image of $\{57,58,\dots,63\}$ under this map evaluated over $\mathbb{F}_{2}^{6}$ with an all-ones row appended. This yields
$\begin{pmatrix}1&1&1&1&1&0&1\\ 1&1&1&1&1&1&0\\ 1&1&0&1&1&1&0\\ 0&0&1&1&1&1&0\\ 0&1&1&0&0&1&1\\ 1&0&1&0&1&0&1\\ 1&1&1&1&1&1&1\end{pmatrix}$
which is invertible. Example code is also provided online [25].
IV Higher-Order Staircase Codes ($L\geq 1$, $M\geq 1$)
Tiled diagonal zipper codes [11] are a family of scattering generalizations of staircase codes for which the encoding and decoding memory approaches half of that of classical staircase codes while maintaining a variable node degree of $M+1=2$. In this section, we further generalize our scattering generalized staircase codes from Section II to both include this family and to extend it to arbitrary bit or variable node degree $M+1$. This generalization has two uses: The first is that the encoding and decoding memory can be reduced while keeping all else essentially constant. The second is that it provides flexibility in hardware designs by allowing the sidelength to be tuned while keeping all else essentially constant. The sidelength is a proxy for the degree of parallelism that is naturally admitted by an encoding or decoding circuit. We refer to the resulting scattering code family as higher-order staircase codes since it subsumes and extends the generalized staircase codes of Definition 7.
IV-A Code Structure
Consider starting with a weak generalized staircase code of sidelength $S^{\prime}$ and degree parameter $M^{\prime}$. We further have our permutations of $[S^{\prime}]\times[S^{\prime}]$ given by $\pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$ and the corresponding matrix-permuting functions given by $\Pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$. Finally, we have a ruler $(d_{0},d_{1},\dots,d_{M^{\prime}})$ and a component code $\mathcal{C}$ of length $(M^{\prime}+1)S^{\prime}$ and dimension $(M^{\prime}+1)S^{\prime}-r$. However, we have the following twist: We define our constraint that the rows of
$\big{(}\Pi^{\prime}_{M^{\prime}}(B_{n-d_{M^{\prime}}})\big{|}\cdots\big{|}\Pi^{\prime}_{1}(B_{n-d_{1}})\big{|}\Pi^{\prime}_{0}(B_{n-d_{0}})\big{)}$ (14)
belong to $\mathcal{C}$ as holding for all integers $n\in L\mathbb{Z}$ where $L$ is a new fixed positive integer parameter. In other words, we restrict the constraint on the rows of (14) to integers $n$ that are multiples of $L$.
Revisiting the Tanner graph perspective, the constraint nodes become $\mathcal{L}=L\mathbb{Z}\times[S^{\prime}]$ and from (5) we see that a variable node $(n,i,j)\in\mathcal{P}$ has the constraint $(n^{\prime},i^{\prime})\in\mathcal{L}$ as a neighbour if and only if $n^{\prime}=n+d_{k}$ for some $k\in[M^{\prime}+1]$ where $n^{\prime}\in L\mathbb{Z}$. Therefore, we get a variable node degree of
$|\absolutevalue{\mathcal{N}((n,i,j))}=\absolutevalue{\{d_{k}\,|\,k\in[M^{\prime}+1],d_{k}\equiv-n\pmod{*}{L}\}}$ | (15)
which depends on the block index $n$ and the ruler. We seek to recover our original regular variable degree of $M+1$ for all variable nodes independently of $n$. This requires an $L$-uniform ruler defined as follows:
Definition 8 ($L$-uniform ruler).
A ruler $(d_{0},d_{1},\dots,d_{M^{\prime}})$ of order $M^{\prime}+1$ is $L$-uniform if it has an equal number of marks belonging to each of the residue classes
$L\mathbb{Z},L\mathbb{Z}+1,\dots,L\mathbb{Z}+L-1\text{.}$
Necessarily, any such ruler can be constructed from a set of $L$ base rulers $\mathcal{X}=\{X_{0},X_{1},\dots,X_{L-1}\}$ each of order $M+1=(M^{\prime}+1)/L$ as the union of the entries of
$LX_{0},LX_{1}+1,\dots,LX_{L-1}+L-1$
together with some ordering.
Remark 4.
Normalizing an $L$-uniform ruler produces a ruler which is still $L$-uniform.
Remark 5.
Given $L$ order $M+1$ (base) rulers
$X_{\ell}=(d_{0}^{(\ell)},d_{1}^{(\ell)},\dots,d_{M}^{(\ell)})$
for $\ell\in[L]$, the elements of the corresponding $L$-uniform ruler of order $L(M+1)$ can be arranged into an $L\times(M+1)$ array
$\begin{array}[]{llll}Ld_{0}^{(0)}&Ld_{1}^{(0)}&\cdots&Ld_{M}^{(0)}\\ Ld_{0}^{(1)}+1&Ld_{1}^{(1)}+1&\cdots&Ld_{M}^{(1)}+1\\ \phantom{L}\vdots&\phantom{L}\vdots&\ddots&\phantom{L}\vdots\\ Ld_{0}^{(L-1)}\!+\!L\!-\!1&Ld_{1}^{(L-1)}\!+\!L\!-\!1&\cdots&Ld_{M}^{(L-1)}\!+\!L\!-\!1\\ \end{array}$
where the rows partition it into parts that are congruent modulo L and the columns partition it into parts that are not congruent modulo L.
From (15) and Definition 8, we see that an $L$-uniform ruler of order $L(M+1)$ leads to
|$\absolutevalue{\mathcal{N}((n,i,j))}=M+1$
for all $(n,i,j)\in\mathcal{P}$ as needed. Therefore, we can henceforth assume an $L$-uniform ruler as in Remark 5. We assume without loss of generality that both the base rulers and the corresponding $L$-uniform ruler are normalized. Before describing the encoding process for the family of codes obtained so far, we derive the scattering conditions which further reveal a simplification: Only $M+1$ distinct permutations are needed, one for each of the columns of the array in Remark 5.
IV-B Scattering Conditions
Observe that the inter-block scattering condition on (9) becomes that for fixed $n^{\prime}_{1},n^{\prime}_{2}\in L\mathbb{Z}$ with $n^{\prime}_{2}>n^{\prime}_{1}$,
$n^{\prime}_{2}-n^{\prime}_{1}=d_{k^{\prime}_{2}}-d_{k^{\prime}_{1}}$
should have a unique solution $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ if any which necessarily satisfies $k_{2}^{\prime}>k^{\prime}_{1}$ and $d_{k^{\prime}_{1}}\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$ since $n^{\prime}_{2}-n^{\prime}_{1}\equiv 0\pmod{*}{L}$. We must then have for some $\ell\in[L]$ that $d_{k^{\prime}_{1}},d_{k^{\prime}_{2}}\in\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1]\}$ thus
$n^{\prime}_{2}-n^{\prime}_{1}=d_{k^{\prime}_{2}}-d_{k^{\prime}_{1}}=L\cdot(d^{(\ell)}_{k_{2}}-d^{(\ell)}_{k_{1}})$ (16)
for some $\ell\in[L]$ and $k_{1},k_{2}\in[M+1]$ with $k_{2}>k_{1}$. Therefore, inter-block scattering is obtained when (16) has a unique solution for $\ell\in[L]$ and $k_{1},k_{2}\in[M+1]$ with $k_{2}>k_{1}$.
Secondly, the intra-block scattering condition on (10) becomes that
$\pi^{\prime}_{k^{\prime}_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi^{\prime}_{k^{\prime}_{2}}(i^{\prime}_{2},j^{\prime}_{2})$ (17)
must have a unique solution for $j^{\prime}_{1},j^{\prime}_{2}\in[S^{\prime}]$. However, the uniqueness of solutions to (17) need only be checked for $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ that solve (16) and thus correspond to $d_{k^{\prime}_{1}}\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$. We can then have $\pi^{\prime}_{k^{\prime}_{1}}=\pi^{\prime}_{k^{\prime}_{2}}$ for all $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ such that $d_{k^{\prime}_{1}}\not\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$. This means that we can repeat every permutation $L$ times so that only $(M^{\prime}+1)/L=M+1$ distinct permutations are needed to get intra-block scattering. In particular, consider a collection of $M+1$ permutations of $[S^{\prime}]\times[S^{\prime}]$ given by $\pi_{k}$ for $k\in[M+1]$ and the corresponding matrix-permuting functions given by $\Pi_{k}$ for $k\in[M+1]$. We can take $\pi^{\prime}_{k^{\prime}}=\pi_{k}$ for every $k\in[M+1]$ and $k^{\prime}\in[M^{\prime}+1]$ such that $d_{k^{\prime}}\in\{Ld^{(\ell)}_{k}+\ell\mid\ell\in[L]\}$. Therefore, given $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$, $k_{1},k_{2}\in[M+1]$, and $\ell\in L$ so that $d_{k^{\prime}_{1}}=Ld^{(\ell)}_{k_{1}}+\ell$ and $d_{k^{\prime}_{2}}=Ld^{(\ell)}_{k_{2}}+\ell$, we have $\pi^{\prime}_{k^{\prime}_{1}}=\pi_{k_{1}}$ and $\pi^{\prime}_{k^{\prime}_{2}}=\pi_{k_{2}}$ so that (17) becomes to
$\pi_{k_{1}}(i^{\prime}_{1},j^{\prime}_{1})=\pi_{k_{2}}(i^{\prime}_{2},j^{\prime}_{2})\text{.}$ (18)
Intra-block scattering then corresponds to uniqueness of solutions to (18) for $j^{\prime}_{1},j^{\prime}_{2}\in[S]$ where $k_{1},k_{2}\in[M+1]$. This is obtained by taking $\pi_{k}$ for $k\in[M+1]$ to correspond to an $(M+1,S^{\prime})$-net.
IV-C Encoding
Note that if the base rulers are normalized, we get that $d_{0}^{(\ell)}=0$ for each $\ell\in[L]$. Observe then from Remark 5 that choosing the natural ordering for the corresponding $L$-uniform ruler so that $d_{0}<d_{1}<\cdots<d_{L(M+1)-1}$ not only results in a normalized ruler, but also results in
$(d_{0},d_{1},\dots,d_{L-1})=(0,1,\dots,L-1)\text{.}$
This leads to
$\pi^{\prime}_{0}=\pi^{\prime}_{1}=\cdots=\pi^{\prime}_{L-1}=\pi_{0}$
where $\pi_{0}$ is taken to be the identity permutation. As a result, the constraint span (14) becomes
$\big{(}\Pi^{\prime}_{M^{\prime}}(B_{n-d_{M^{\prime}}})\big{|}\cdots\big{|}\Pi^{\prime}_{L}(B_{n-d_{L}})\big{|}B_{n-(L-1)}\big{|}\cdots\big{|}B_{n-1}\big{|}B_{n}\big{)}$
where $n\in L\mathbb{Z}$. Encoding is performed $L$ blocks at a time with the $(n/L)$th $S^{\prime}\times(LS^{\prime}-r)$ block of information bits $B_{\frac{n}{L}}^{\mathsf{info}}$ adjoined to $M^{\prime}-L+1$ permuted previous blocks to form
$\big{(}\Pi^{\prime}_{M^{\prime}}(B_{n-d_{M^{\prime}}})\big{|}\cdots\big{|}\Pi^{\prime}_{L+1}(B_{n-d_{L+1}})\big{|}\Pi^{\prime}_{L}(B_{n-d_{L}})\big{|}B_{\frac{n}{L}}^{\mathsf{info}}\big{)}$
whose rows are encoded with systematic encoders for the component code $\mathcal{C}$ which is of dimension $(M^{\prime}+1)S^{\prime}-r$ and length $(M^{\prime}+1)S^{\prime}$. This produces an $S^{\prime}\times r$ by block of parity bits $B_{\frac{n}{L}}^{\mathsf{parity}}$ which is adjoined to the information block to produce $L$ consecutive blocks as
$\big{(}B_{n-(L-1)}\big{|}\cdots\big{|}B_{n-1}\big{|}B_{n}\big{)}=\big{(}\underbrace{\text{---}\;B_{\frac{n}{L}}^{\mathsf{info}}\;\text{---}}_{S^{\prime}\times(LS^{\prime}-r)}\;\big{|}\;\underbrace{\text{--}\;B_{\frac{n}{L}}^{\mathsf{parity}}\;\text{--}}_{S^{\prime}\times r}\big{)}$
[…]
To maintain the same code rate and component code $\mathcal{C}$ independently of $L$, we take
$S^{\prime}=\frac{S}{L}$
so that the component code length is $S^{\prime}(M^{\prime}+1)=S(M+1)$ and the code rate is
$R_{\mathsf{unterminated}}=1-\frac{r}{LS^{\prime}}=1-\frac{r}{S}\text{.}$
IV-D Higher-Order Staircase Codes
It remains to define difference triangle sets (DTSs):
Definition 9 ($(L,M)$-DTS [6, Part VI: Ch. 19]).
An $(L,M)$-DTS is set of $L$ Golomb rulers $\mathcal{X}=\{X_{0},X_{1},\dots,X_{L-1}\}$ each of order $M+1$ whose respective sets of distances are disjoint. Equivalently, $L$ order $M+1$ rulers given by
$X_{\ell}=(d_{0}^{(\ell)},d_{1}^{(\ell)},\dots,d_{M}^{(\ell)})$
where $\ell\in[L]$ form an $(L,M)$-DTS if all differences $d^{(\ell)}_{k_{1}}-d^{(\ell)}_{k_{2}}$ for $k_{1},k_{2}\in[M+1]$ with $k_{1}\neq k_{2}$ and $\ell\in[L]$, are distinct.
It then follows that an $L$-uniform ruler of order $M^{\prime}+1=L(M+1)$ constructed by taking the base rulers to be an $(L,M)$-DTS leads to unique solutions to (16) thus inter-block scattering. The construction is summarized by the following definition:
Definition 10 (Higher-order staircase code).
Given all three of the following:
  • an $(L,M)$-DTS given by
    $0=d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
    for $\ell\in[L]$ with corresponding $L$-uniform ruler of order $L(M+1)$
    $d_{0}<d_{1}<\cdots<d_{L(M+1)-1}$
    given accordingly as
    $\{d_{k}\mid k\in[L(M+1)]\}=\\ \{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1],\ell\in[L]\}\text{;}$
  • an $(M+1,S/L)$-net with corresponding $M+1$ permutations of $[S/L]\times[S/L]$ given by $\pi_{k}$ for $k\in[M+1]$ where $\pi_{0}$ is the identity permutation, and a resulting collection of $L(M+1)$ permutations of $[S/L]\times[S/L]$ given by $\pi^{\prime}_{k^{\prime}}=\pi_{k}$ for every $k\in[M+1]$ and $k^{\prime}\in[L(M+1)]$ such that $d_{k^{\prime}}\in\{Ld^{(\ell)}_{k}+\ell\mid\ell\in[L]\}$; and
  • a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$,
a higher-order staircase code of rate $1-r/S$ is defined by the constraint on the bi-infinite sequence of $(S/L)\times(S/L)$ matrices $\dots,B_{-2},B_{-1},B_{0},B_{1},B_{2},\dots$ that the rows of
$\big{(}\Pi^{\prime}_{L(M+1)-1}(B_{n-d_{L(M+1)-1}})\big{|}\cdots\big{|}\Pi^{\prime}_{1}(B_{n-d_{1}})\big{|}\Pi^{\prime}_{0}(B_{n-d_{0}})\big{)}$
belong to $\mathcal{C}$ for all $n\in L\mathbb{Z}$.
Importantly, Section IV-B implies the following:
Proposition 3.
Higher-order staircase codes are scattering.
We provide a concrete example:
Example 4.
Consider taking $L=2$ and $M=2$. The $L=2$ Golomb rulers of order $M+1=3$ given by
$(d_{0}^{(0)},d_{1}^{(0)},d_{2}^{(0)})$ $=(0,6,7)$
$(d_{0}^{(1)},d_{1}^{(1)},d_{2}^{(1)})$ $=(0,2,5)$
form a $(2,2)$-DTS since their respective sets of distances $\{1,6,7\}$ and $\{2,3,5\}$ are disjoint. We further have
$2(0,6,7)$ $=(0,12,14)$
$2(0,2,5)+1$ $=(1,5,11)$
thus a corresponding $2$-uniform ruler given by
$(d_{0},d_{1},d_{2},d_{3},d_{4},d_{6})=(0,1,5,11,12,14)\text{.}$
Finally, we have
$(2d_{0}^{(0)},2d_{0}^{(1)}+1)$ $=(0,1)=(d_{0},d_{1})$
$(2d_{1}^{(0)},2d_{1}^{(1)}+1)$ $=(12,5)=(d_{4},d_{2})$
$(2d_{2}^{(0)},2d_{2}^{(1)}+1)$ $=(14,11)=(d_{5},d_{3})$
thus
$\pi^{\prime}_{0}$ $=\pi^{\prime}_{1}=\pi_{0}$
$\pi^{\prime}_{4}$ $=\pi^{\prime}_{2}=\pi_{1}$
$\pi^{\prime}_{5}$ $=\pi^{\prime}_{3}=\pi_{2}\text{.}$
We consider the $(3,S/2)$-net permutations of Example 1 which are given by $\pi_{0}(i,j)=(i,j)$, $\pi_{1}(i,j)=(j,i)$, and $\pi_{2}(i,j)=(j,i+j)\pmod{S/2}$. This leads to $\Pi_{0}(B)=B$, $\Pi_{1}(B)=B^{\mathsf{T}}$, and we further define $B^{\pi}=\Pi_{2}(B)$. Our higher-order staircase code is then defined by the constraint that the rows of
$\big{(}B_{n-14}^{\pi}\big{|}B_{n-12}^{\mathsf{T}}\big{|}B_{n-11}^{\pi}\big{|}B_{n-5}^{\mathsf{T}}\big{|}B_{n-1}\big{|}B_{n}\big{)}$
belong to some specified component code $\mathcal{C}$ for all $n\in 2\mathbb{Z}$.
Figure 4: Visualization of a tiled diagonal zipper code corresponding to $L=2$, $M=1$, $(d_{1}^{(0)},d_{1}^{(1)})=(1,2)$, and $\Pi_{1}(B)=B^{\mathsf{T}}$; rows belong to $\mathcal{C}$.
Figure 5: Visualization of the code of Example 4 where $L=2$ and $M=2$; the respective sets of distances of the two Golomb rulers are disjoint and rows belong to $\mathcal{C}$.
A visualization of an $(L=2,M=1)$ code corresponding to an instance of tiled diagonal zipper codes [11] is provided in Fig. 4. A visualization of the $(L=2,M=2)$ code of Example 4 follows it in Fig. 5 illustrating an increase in variable degree while maintaining scattering. In both of these pictures, the $L=2$ rightmost columns represent the blocks that are actually transmitted while the columns to their left are delayed and permuted copies that are used to constrain the transmitted blocks.
Finally, we note that for higher-order staircase codes as just defined, the code rate, $4$-cycle-freeness, component code, and variable and constraint node degrees are all independent of $L$. It is therefore natural to study the effect of increasing $L$ while keeping all else fixed. We will find that while increasing $L$ reduces the block size to $1/L^{2}$ times its size, the number of blocks that must be kept in memory is $\Omega(L^{2})$ meaning that we can at most reduce the memory by a constant factor. However, this factor can still be meaningful in practice and we may nonetheless need to minimize memory for any given choice of $L$. Sections VI and VII will deal with exhibiting DTSs that lead to memory-optimal realizations of higher-order staircase codes.
Before proceeding, we comment briefly on the practical code performance implications of increasing $L$. Since the code rate, degree distribution, $4$-cycle-freeness, and component code are all unaffected by $L$, coding folklore suggests that the threshold and error floor performance should be roughly independent of $L$. For example, in the case of tiled diagonal zipper codes corresponding to $(L\geq 1,M=1)$, it is found in [3] that the gap to the hard-decision Shannon limit is increased by no more than $0.04$ dB as $L$ is increased from $L=1$ to $L=S=1000$ with the component code being a shortened triple-error-correcting BCH code of length $2000$ and dimension $1967$. A similar invariance can be observed for general higher-order staircase codes by experimenting with our simulators for these codes available available online [21, 25]. The reader is also referred to [26] for recent work on theoretical performance analysis under similar code families and decoding methods to what is considered in this paper. Moreover, the bound of Proposition 1 is independent of $L$ providing the same guarantee on the coarse asymptotic behaviour of the error floor.
IV-E Multiply-Chained and Other Extensions
We now adapt a technique of Zhao et al. [13] in which multiple tiled diagonal zipper codes [11] are chained in a circle to boost their performance. This is straightforwardly applicable to higher-order staircase codes. We introduce a new positive integer parameter $C>1$ and consider $C$ higher-order staircase code encoders indexed by $c\in[C]$ and acting on $C$ independent input streams with the following twist: The $c$th encoder uses the memory of the $(c-1\pmod{*}{C})$th encoder. This leads to multiply-chained higher-order staircase codes formally defined as follows:
Definition 11 (Multiply-chained higher-order staircase code).
Given all four of the following:
  • an integer $C>1$;
  • an $(L,M)$-DTS given by
    $0=d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
    for $\ell\in[L]$ with corresponding $L$-uniform ruler of order $L(M+1)$
    $d_{0}<d_{1}<\cdots<d_{L(M+1)-1}$
    given accordingly as
    $\{d_{k}\mid k\in[L(M+1)]\}=\\ \{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1],\ell\in[L]\}\text{;}$
  • an $(M+1,S/L)$-net with corresponding $M+1$ permutations of $[S/L]\times[S/L]$ given by $\pi_{k}$ for $k\in[M+1]$ where $\pi_{0}$ is the identity permutation, and a resulting collection of $L(M+1)$ permutations of $[S/L]\times[S/L]$ given by $\pi^{\prime}_{k^{\prime}}=\pi_{k}$ for every $k\in[M+1]$ and $k^{\prime}\in[L(M+1)]$ such that $d_{k^{\prime}}\in\{Ld^{(\ell)}_{k}+\ell\mid\ell\in[L]\}$; and
  • a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$,
a multiply-chained higher-order staircase code of rate $1-r/S$ is defined by the constraint on the $C$ bi-infinite sequences of $(S/L)\times(S/L)$ matrices $\dots,B_{-2}^{(c)},B_{-1}^{(c)},B_{0}^{(c)},B_{1}^{(c)},B_{2}^{(c)},\dots$ indexed by $c\in[C]$ that the rows of
$\bigg{(}\Pi^{\prime}_{L(M+1)-1}(B_{n-d_{L(M+1)-1}}^{(c-1\pmod{*}{C})})\bigg{|}\\ \cdots\bigg{|}\Pi^{\prime}_{L}(B_{n-d_{L}}^{(c-1\pmod{*}{C})})\bigg{|}\Pi^{\prime}_{L-1}(B_{n-d_{L-1}}^{(c)})\bigg{|}\cdots\bigg{|}\Pi^{\prime}_{0}(B_{n-d_{0}}^{(c)})\bigg{)}$
for each $c\in[C]$ belong to $\mathcal{C}$ for all $n\in L\mathbb{Z}$.
We define a multiply-chained higher-order staircase code with $C=1$ to simply be a higher-order staircase code as in Definition 10 so that we have a unified family defined for all $C\geq 1$. Lastly, we include one further generalization of this family which is as follows: Rather than considering one component code $\mathcal{C}$, we can consider a “time-varying” family of component codes such as $\{\mathcal{C}_{i}^{(c)}\mid c\in[C],i\in[S/L]\}$ where the $i$th row in the constraint span in Definition 11 belongs to $\mathcal{C}_{i}^{(c)}$. For example, this family can simply consist of $S/L$ permuted copies of a fixed code $\mathcal{C}$. With this extension, the family of Definition 11 includes the OFEC code family of [12] as the special case of $C=2$, $M=1$, and a special choice of time-varying component code permutation defined therein.
IV-F Memory
We now consider the encoding memory requirements for higher-order staircase codes. For the multiply-chained variation, all requirements are simply multiplied by $C$ so we restrict our attention to the case of $C=1$. At first glance, encoding requires that memory of the last $d_{M^{\prime}}$ transmitted blocks be maintained. A naive choice would be to maintain the set of $L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}$ previous blocks $\mathsf{BUFF}_{\frac{n}{L}}$ defined by
$\mathsf{BUFF}_{\frac{n}{L}}=\bigg{\{}B_{n-\mu}\;\bigg{|}\;\\ \mu\in\left\{L,L+1,\dots,L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}+L-1\right\}\bigg{\}}$
for encoding the $(n/L)$th information block. However, we can observe in the case of Example 4 illustrated in Fig. 5 that blocks with even indices must be recalled for longer than blocks with odd indices. In general, by examining the ruler decomposition given in Remark 5, we see that we can alternatively maintain $L$ independent buffers $\mathsf{BUFF}^{(\ell)}_{\frac{n}{L}}$ of sizes $d_{M}^{(\ell)}$ for $\ell\in[L]$ given by
$\mathsf{BUFF}^{(\ell)}_{\frac{n}{L}}=\{B_{n-\mu}\mid\mu=L\nu+\ell,\nu\in\{1,2,\dots,d_{M}^{(\ell)}\}\}$
so that the total number of blocks stored is
$\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}\leq L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}\text{.}$
We then have an encoding memory of
$\left(\frac{S}{L}\right)^{2}\cdot\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ (19)
bits (or nonbinary symbols). This is analogous to the concept of total memory or overall constraint length in the theory of convolutional codes [27, Ch. 2] with the key differences being that bits are replaced with $(S/L)\times(S/L)$ blocks which admit permutation operations and that one-bit memory elements are replaced accordingly with stores of $(S/L)^{2}$ bits. The encoder style assumed here is analogous to Forney’s obvious realization [28].
On the other hand, the decoding window must typically be chosen as a number of blocks which is a constant multiple of $d_{L(M+1)-1}+1$, the smallest number of consecutive blocks which contains a complete component codeword. If we order our base rulers in descending order of length so that
$\max_{\ell\in[L]}d_{M}^{(\ell)}=d_{M}^{(0)}>d_{M}^{(1)}>\cdots>d_{M}^{(L-1)}$
thus $d_{L(M+1)-1}=L\cdot d_{M}^{(0)}=L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}$, we get
$\left(\frac{S}{L}\right)^{2}\cdot\left(1+L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}\right)$ (20)
as a measure of the decoding memory. This is analogous to the constraint length in the theory of convolutional codes.
We emphasize that (19) and (20) are merely models for the encoding and decoding memory under a certain encoder and decoder structure considered in [11] which can be subverted in practice. There are two main ways in which this is done:
  • Each $(S/L)\times(S/L)$ block is compressed down to a $(S/L)\times r$ future additive contribution to parity and thus is not directly stored.
  • An encoder style analogous to the observer canonical form as opposed to the controller canonical form [27, Ch. 2] or obvious realization [28] is used in which the encoding memory is proportional to $\max_{\ell\in[L]}d_{M}^{(\ell)}$ rather than $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$. This encoder style was first recognized by Massey [29].
Both of these are done in the hardware implementations of a staircase code encoder corresponding to $(L=1,M=1)$ and a zipper code encoder corresponding to $(L=32,M=1)$ found in [30] and [31] respectively. It is found in [31] that there are significant hardware efficiency benefits to $L=32$ relative to $L=1$. Nonetheless, one or both of $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ and $\max_{\ell\in[L]}d_{M}^{(\ell)}$ are measures of hardware complexity. Thus, we seek, for a given $L$ and $M$, an $(L,M)$-DTS for which $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ and/or $\max_{\ell\in[L]}d_{M}^{(\ell)}$ are minimized. We defer the study of this problem to Section VI.
V Extended-Hamming-Based Higher-Order Staircase Codes
We consider now the use of memory-optimal higher-order staircase codes and their multiply-chained variation with extended Hamming component codes. As discussed in the previous section, the construction of such codes requires the construction of DTSs with minimum scope and sum-of-lengths. Such DTSs are provided in the Appendix with the subject of how they are obtained deferred to Sections VI and VII. We use the optimal DTSs from the Appendix together with the extended Hamming codes defined by Table V to construct our codes.
The encoder produces a $C(S/L)\times L(S/L)$ rectangle of coded bits since each of the $C$ higher-order staircase code encoders produces $L$ blocks at a time each of dimensions $(S/L)\times(S/L)$. The decoding window size $W$ counts the number of such rectangles meaning that the number of bits in the decoding window is $WC(S/L)^{2}L$. Similarly to before, we assume transmission across a BSC and perform sliding window decoding of $W$ consecutive rectangles at a time where one iteration comprises decoding all constraints in the window consecutively and $I$ iterations are performed before advancing the window by one rectangle. We again use a termination-like procedure as in Section III in which a frame is formed from $F$ consecutive rectangles. We take $F=10^{5}+W$ for all simulations in this section so that the entire coding scheme is determined by the six-tuple of parameters $(L,M,S/L,C,W,I)$. The simulation curves in Figures 6 and 7 are labeled with this six-tuple appended with the resulting unterminated code rate $R_{\mathsf{unterminated}}$ and the decoding window size in bits for the reader’s convenience.
Figure 6: Simulation results for $R_{\mathsf{unterminated}}\approx 0.937$ examples with parameters $(L,M,S/L,C,W,I,R_{\mathsf{unterminated}},WC(S/L)^{2}L)$
Figure 7: Simulation results for $R_{\mathsf{unterminated}}\approx 0.87$ examples with parameters $(L,M,S/L,C,W,I,R_{\mathsf{unterminated}},WC(S/L)^{2}L)$
One of our goals in this section is to demonstrate the ability to construct a variety of codes with similar properties by varying $L$ at a roughly constant code rate. We accordingly organize our simulations slightly differently with results for two different approximate code rates given in Figures 6 and 7 respectively. The dashed lines indicate points at which zero bit errors were measured after transmitting a number of bits equal to ten times the reciprocal of the corresponding output BER.
Observe firstly that Fig. 7 illustrates control over the error floor and threshold via $M$ and $C$ respectively as predicted. Moreover, the examples in these figures perform comparably to the codes compared in [2, 13] while using only Hamming components as opposed to BCH components and also improve over the results of Section III with generalized staircase codes $(L=1)$ in terms of memory and decoding iterations.
We provide heuristic estimates of the performance, latency, and complexity for some examples as before. In Table VII, we compare the classical staircase code of [2] to the extended-Hamming-based higher-order staircase code with parameters $(L,M,S/L,C,W,I)=(7,4,25,1,162,1)$. As before, we assume that for the relatively small values of $M$ considered, the cost of each component code decoding is dominated by the cost of invoking the component code decoder $t^{2}$ rather than the cost of updating the syndromes of codewords involved in each corrected bit. In Table VIII, we compare a multiply-chained tiled diagonal zipper code from [13] with parameters $(L,M,S/L,C,W,I)=(2,1,125,2,8,4)$ and BCH component codes with $t=3$ to the extended-Hamming-based higher-order staircase code with parameters $(L,M,S/L,C,W,I)=(4,4,19,1,96,1)$. In both situations, we see large reductions in latency and complexity at a performance cost of roughly $0.3$ to $0.5$ dB.
TABLE VII: Heuristic estimates of performance, latency, and complexity for higher-order staircase codes with $R_{\mathsf{unterminated}}\approx 0.937$
staircase [2, 22] this work
gap to Shannon at 1e-15 0.56 dB 0.89 dB
latency: $WC(S/L)^{2}L$ 1.8e6 7.1e5
iterations: $I$ 3 1
decodings per iteration: $WC(S/L)$ 3.6e3 4.1e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWC(S/L)t^{2}$ 9.7e4 4.1e3
TABLE VIII: Heuristic estimates of performance, latency, and complexity for higher-order staircase codes with $R_{\mathsf{unterminated}}\approx 0.87$
Fig. 13 in [13] this work
gap to Shannon at 1e-8 0.75 dB 1.25 dB
latency: $WC(S/L)^{2}L$ 5e5 1.4e5
iterations: $I$ 4 1
decodings per iteration: $WC(S/L)$ 2e3 1.8e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWC(S/L)t^{2}$ 7.2e4 1.8e3
We anticipate improved performance–complexity–latency tradeoffs in a wider variety of contexts in light of recent works [32, 33] which accomplish such goals via new product-like codes with bit degree increased from two to three. The simulations here are only meant to provide a few motivating examples. These results can be reproduced and extended using our simulation tools available online [21, 25].
VI Difference Triangle Sets with Minimum Scope and Sum-of-Lengths
Consider an $(L,M)$-DTS given by
$0=d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
for $\ell\in[L]$. As with Golomb rulers or $(1,M)$-DTSs, it is easy to construct an $(L,M)$-DTS for arbitrary $L$ and $M$ if no further objectives are placed. A well-studied hard problem is that of constructing an $(L,M)$-DTS of minimum scope where scope is the largest mark across all (normalized) rulers $\max_{\ell\in[L]}d_{M}^{(\ell)}$. Equivalently, scope is the maximum length among all rulers. We again have a trivial bound on the scope of an $(L,M)$-DTS given by
$\max_{\ell\in[L]}d_{M}^{(\ell)}\geq L\cdot\frac{(M+1)M}{2}$ (21)
with equality if and only if the union of the sets of distances of the $L$ rulers is precisely $\{1,2,\dots,L(M+1)M/2\}$. An $(L,M)$-DTS for which the trivial bound holds with equality is said to be perfect. The union of the $L$ (disjoint) sets of distances of the $L$ rulers forming an $(L,M)$-DTS will be referred to concisely as its set of distances.
On the other hand, we also require a minimization of the sum-of-lengths $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$. To the best of the authors’ knowledge, this problem has not been explicitly studied. We begin by identifying when minimizing scope is equivalent to minimizing sum-of-lengths. Aside from the case of $L=1$ where sum-of-lengths and scope are trivially equivalent, we have the following:
Proposition 4.
A minimum scope $(L,M)$-DTS also has minimum sum-of-lengths if $M\in\{1,2\}$ and it is perfect.
Proof.
In the case of $M=2$, observe that
$2\sum_{\ell=0}^{L-1}d_{2}^{(\ell)}$ $=\sum_{\ell=0}^{L-1}((d_{2}^{(\ell)}-d_{1}^{(\ell)})+d_{1}^{(\ell)}+d_{2}^{(\ell)})$ (22)
$\geq\frac{3L(3L+1)}{2}\text{.}$ (23)
In particular, (22) is a sum over all three distances of all $L$ rulers which must be distinct positive integers. This leads to the lower bound (23) which is achieved precisely when perfect. When $M=1$, the sum of the lengths is precisely the sum of the distances which means that a similar argument holds. ∎
Beyond this case, the problem of minimizing sum-of-lengths will be distinct and non-trivial. If $M>2$, there is no simple proportionality relationship such as (22) between a sum over all ruler lengths and a sum over all distances. Moreover, even in the case of $M=2$, a minimum scope $(L,2)$-DTS which is not perfect does not necessarily minimize sum-of-lengths as illustrated by the following counterexample:
Example 5.
Consider the $(2,2)$-DTS of Example 4 given by $\{(0,2,5),(0,6,7)\}$ which is of scope $7$. Observe that this scope is already minimum since achieving the scope lower bound of $6$ is impossible since it requires equality of (22) which is even and (23) which is $21$. However, the alternative $(2,2)$-DTS given by $\{(0,3,4),(0,2,7)\}$ which also has minimum scope also has a smaller sum-of-lengths.
Except for the cases of $M=1$ and $M=2$, it remains an open problem to exhibit minimum scope $(L,M)$-DTSs for arbitrary $L$. However, perfect DTSs admit recursive constructions which lead to constructive existence results for infinitely many values of $L$. It is known that perfect DTSs can only exist for $M\leq 4$. Therefore, for the cases of $M=3$ and $M=4$, recursive constructions provide the strongest existence results. The values of $L$ covered by such constructions are often (but not always) too large to be of relevance to applications. Therefore, such constructions do not eliminate the need for computer-search constructions for smaller values of $L$ but are nonetheless at least of theoretical interest. The reader is referred to [34, 35] and the references therein for more on this context. We will study the implications of one such construction due to Wild [36] and Kotzig and Turgeon [37] for the problem of minimizing sum-of-lengths. We will need the following:
Definition 12 (Sharply $2$-transitive permutation group).
A sharply $2$-transitive permutation group $G$ acting on a set $X$ is a group of permutations of $X$ with the following property: For each pair of ordered pairs $(w,x),(y,z)\in X\times X$ with $w\neq x$ and $y\neq z$, there exists a unique permutation $g\in G$ such that $(g(w),g(x))=(y,z)$.
As noted in [36], we can exhibit a finite such group as the group of invertible affine transformations from a finite field $\mathbb{F}_{M+1}$ to itself $x\mapsto ax+b$ where $M+1$ is a prime power and $a,b\in\mathbb{F}_{M+1}$ with $a\neq 0$ yielding $M(M+1)$ such permutations. For $M=3$, the resulting permutations of $[M+1]$ in array format are
$\begin{array}[]{cccccccccccc}0&1&2&3&0&1&2&3&0&1&2&3\\ 1&0&3&2&2&3&0&1&3&2&1&0\\ 2&3&0&1&3&2&1&0&1&0&3&2\\ 3&2&1&0&1&0&3&2&2&3&0&1\\ \end{array}$
and for $M=4$, the permutations are defined by simple modular arithmetic since $\mathbb{F}_{5}$ is $\mathbb{Z}_{5}$.
Theorem 5 (Special case of the combining construction of Wild [36] and Kotzig and Turgeon [37]).
Let $M+1$ be a prime power and let the collection of $M(M+1)$ permutations
$\rho_{k}\colon[M+1]$ $\longrightarrow[M+1]$
$i$ $\longmapsto\rho_{k}(i)$
where $k\in[M(M+1)]$ be a sharply $2$-transitive permutation group acting on $[M+1]$. Let $\mathcal{X}$ be a perfect $(L_{1},M)$-DTS and $\mathcal{Y}$ be a perfect $(L_{2},M)$-DTS with each given respectively by
$\mathcal{X}=\{(x_{0}^{(i)},x_{1}^{(i)},\dots,x_{M}^{(i)})\mid i\in[L_{1}]\}$
and
$\mathcal{Y}=\{(y_{0}^{(i)},y_{1}^{(i)},\dots,y_{M}^{(i)})\mid i\in[L_{2}]\}\text{.}$
Let $\mathcal{W}$ be the set of $L_{1}L_{2}M(M+1)$ rulers given by
$(L_{2}M(M+1)+1)\left(x_{0}^{(i)},x_{1}^{(i)},\dots,x_{M}^{(i)}\right)+\\ \left(y_{\rho_{k}(0)}^{(j)},y_{\rho_{k}(1)}^{(j)},\dots,y_{\rho_{k}(M)}^{(j)}\right)$
for $i\in[L_{1}]$, $j\in[L_{2}]$, and $k\in[M(M+1)]$. Let $\mathcal{Z}$ be the union of $\mathcal{Y}$, $(L_{2}M(M+1)+1)\mathcal{X}$, and $\mathcal{W}$. Then, $\mathcal{Z}$ is a perfect $(L_{1}L_{2}M(M+1)+L_{1}+L_{2},M)$-DTS. (However, the rulers are in unnormalized form.)
Proposition 6 (Sum-of-lengths under combining).
Combining a perfect $(L_{1},M)$-DTS with sum-of-lengths $S_{1}$ and a perfect $(L_{2},M)$-DTS with sum-of-lengths $S_{2}$ as in Theorem 5 yields a perfect $(L_{1}L_{2}M(M+1)+L_{1}+L_{2},M)$-DTS with sum-of-lengths $(L_{2}M(M+1)+1)^{2}S_{1}+S_{2}$.
Proof.
Trivially, $\mathcal{Y}$ and $(L_{2}M(M+1)+1)\mathcal{X}$ contribute $S_{2}$ and $(L_{2}M(M+1)+1)S_{1}$ to the sum-of-lengths so it remains to calculate the contribution of $\mathcal{W}$. The difference between the $v$th and $u$th mark of the ruler in $\mathcal{W}$ corresponding to a fixed $i\in[L_{1}]$, $j\in[L_{2}]$, and $k\in[M(M+1)]$ can be expressed as
$(L_{2}M(M+1)+1)\left(x_{v}^{(i)}-x_{u}^{(i)}\right)+\left(y_{\rho_{k}(v)}^{(j)}-y_{\rho_{k}(u)}^{(j)}\right)\text{.}$
Since the term $y_{\rho_{k}(v)}^{(j)}-y_{\rho_{k}(u)}^{(j)}$ satisfies
$-\frac{L_{2}M(M+1)}{2}\leq y_{\rho_{k}(v)}^{(j)}-y_{\rho_{k}(u)}^{(j)}\leq\frac{L_{2}M(M+1)}{2}\text{,}$
the scale factor $(L_{2}M(M+1)+1)$ guarantees that the ordering of the differences $x_{v}^{(i)}-x_{u}^{(i)}$ is preserved. Assuming without loss of generality that the rulers of $\mathcal{X}$ are normalized so that the length of the $i$th ruler of $\mathcal{X}$ is the difference corresponding to $v=M$ and $u=0$, the same difference will then yield the length of the corresponding rulers in $\mathcal{W}$ as
$(L_{2}M(M+1)+1)x_{M}^{(i)}+\left(y_{\rho_{k}(M)}^{(j)}-y_{\rho_{k}(0)}^{(j)}\right)\text{.}$ (24)
Lastly, sharp $2$-transitivity guarantees that
$\sum_{k=0}^{M(M+1)-1}y_{\rho_{k}(M)}^{(j)}-y_{\rho_{k}(0)}^{(j)}=0$
in which case summing (24) over all $i\in[L_{1}]$, $j\in[L_{2}]$, and $k\in[M(M+1)]$, yields
$L_{2}M(M+1)(L_{2}M(M+1)+1)S_{1}$
which is added to $S_{2}$ and $(L_{2}M(M+1)+1)S_{1}$ to obtain the desired result. ∎
We will now consider each of the cases of $M\in\{1,2,3,4\}$. These cases are both practically the most relevant for applications to higher-order staircase codes but are also theoretically interesting for the reasons just discussed.
VI-A The Case of $M=1$
For $(L,1)$-DTSs, we have the trivial bound
$\max_{\ell\in[L]}d_{1}^{(\ell)}\geq L$
and trivially
$\sum_{\ell=0}^{L-1}d_{1}^{(\ell)}\geq\frac{L(L+1)}{2}$ (25)
with both of these bounds holding with equality for the perfect $(L,1)$-DTS given by $\{(0,1),(0,2),\dots,(0,L)\}$.
VI-B The Case of $M=2$
For $(L,2)$-DTSs, it is known that comparing the parities of (22) and (23) slightly strengthens the trivial bound to
$\max_{\ell\in[L]}d_{2}^{(\ell)}\geq\begin{cases}3L&\text{if }L\equiv 0\text{ or }1\pmod{*}{4}\\ 3L+1&\text{if }L\equiv 2\text{ or }3\pmod{*}{4}\end{cases}\text{.}$ (26)
From this and (22), it straightforwardly follows that
$\sum_{\ell=0}^{L-1}d_{2}^{(\ell)}\\ \geq\begin{dcases}\frac{3L(3L+1)}{4}&\text{if }L\equiv 0\text{ or }1\pmod{*}{4}\\ \frac{(3L-1)3L}{4}+\frac{3L+1}{2}&\text{if }L\equiv 2\text{ or }3\pmod{*}{4}\end{dcases}$ (27)
where the first inequality holds with equality if and only if the set of distances is precisely $\{1,2,\dots,3L\}$ and the second inequality holds with equality if and only if set of distances is precisely $\{1,2,\dots,3L-1,3L+1\}$.
An explicit construction of $(L,2)$-DTSs achieving the bound (26) exists for arbitrary $L$ and was given by Skolem [38] in the case of $L\equiv 0\text{ or }1\pmod{*}{4}$ and by O’Keefe [39] in the case of $L\equiv 2\text{ or }3\pmod{*}{4}$. In the case of $L\equiv 0\text{ or }1\pmod{*}{4}$, by perfectness, the bound (27) is automatically achieved. In the case of $L\equiv 2\text{ or }3\pmod{*}{4}$, the construction of [39] happens to preclude $3L$ from the set of distances and hence also achieves (27). Note that Example 5 demonstrates how it is possible to achieve (26) without achieving (27). Nonetheless, the specific constructions of [38] and [39] simultaneously achieve (26) and (27) for arbitrary $L$ settling this case. We provide the constructions of [38, 39] in the Appendix for the reader’s convenience since [38, 39] consider a different but equivalent problem to the $(L,2)$-DTS problem requiring a well-known but slightly tedious translation.
VI-C The Case of $M=3$
For $(L,3)$-DTSs, the trivial bound is
$\max_{\ell\in[L]}d_{3}^{(\ell)}\geq 6L$ (28)
and it is a longstanding conjecture of Bermond [40] that this is achievable for all $L$ excluding $L\in\{2,3\}$. Theorem 5 implies that there are infinitely many values of $L$ for which (28) is achieved, by, for example, repeatedly combining the perfect $(1,3)$-DTS $\{(0,1,4,6)\}$ with itself. Bermond’s conjecture has been verified for all $L\leq 1000$ in [41] by using a wide variety of recursive constructions to cover gaps as needed.
Proposition 7.
The sum-of-lengths of an $(L,3)$-DTS satisfies
$\sum_{\ell=0}^{L-1}d_{3}^{(\ell)}\geq 5L^{2}+L\text{.}$ (29)
Proof.
By the distinctness of positive differences, we have
$\sum_{u=1}^{6L}u$ $\leq\sum_{\ell=0}^{L-1}\sum_{i=0}^{3}\sum_{j=i+1}^{3}d_{j}^{(\ell)}-d_{i}^{(\ell)}$
$=4\sum_{\ell=0}^{L-1}d_{3}^{(\ell)}-\sum_{\ell=0}^{L-1}(d^{(\ell)}_{1}-d^{(\ell)}_{0})+(d^{(\ell)}_{3}-d^{(\ell)}_{2})$
$\leq 4\sum_{\ell=0}^{L-1}d_{3}^{(\ell)}-\sum_{u=1}^{2L}u$
which yields the desired result after rearranging. ∎
It is not the case that $(L,3)$-DTSs achieving (28) such as those listed in [42] achieve (29). However, we conjecture that it is possible to simultaneously achieve (28) and (29) for all $L$ excluding $L\in\{2,3,4,5\}$. By using a computer-search technique described in Section VII, we found $(L,3)$-DTSs simultaneously achieving (28) and (29) for all $6\leq L\leq 15$ and these are provided in the Appendix. In the cases of $L\in\{2,3,4,5\}$, one or both of (28) and (29) are not achievable but simultaneous absolute minimization of scope and sum-of-lengths is still possible and the optimal solutions are provided in the Appendix. These cases are small enough to be handled by an exact mixed-integer linear programming (MILP) solver by slightly modifying the MILP formulation of the minimum scope DTS problem given in [43]. Alternatively, an exhaustive backtracking DTS search program of Shearer [44, 45] can be used for these small cases.
Next, we consider constructing an infinite family via Theorem 5. For all integers $i\geq 0$, let $\mathcal{Z}_{i}$ be a perfect $(L_{i},3)$-DTS with sum-of-lengths $S_{i}$ obtained by combining $\mathcal{Z}_{i-1}$ and $\mathcal{Z}_{0}$ via Theorem 5 where $\mathcal{Z}_{0}$ is a fixed perfect $(L_{0},3)$-DTS with sum-of-lengths $S_{0}$. Proposition 6 provides linear recurrences for $L_{i}$ and $S_{i}$ which we can solve to get
$L_{i}=\frac{1}{12}\left((12L_{0}+1)^{i+1}-1\right)$
and
$S_{i}=\frac{6S_{0}}{6L_{0}^{2}+L_{0}}L_{i}^{2}+\frac{S_{0}}{6L_{0}^{2}+L_{0}}L_{i}\text{.}$
[…]
$S_{i}=\left(5+\frac{1}{6L_{0}+1}\right)L_{i}^{2}+\left(1-\frac{L_{0}}{6L_{0}+1}\right)L_{i}\text{.}$
Since the largest $L_{0}$ for which we have exhibited a perfect $(L_{0},3)$-DTS achieving (29) in the Appendix is $L_{0}=15$, the best result we can get is as follows:
Proposition 8.
There are infinitely many values of $L$ for which a perfect $(L,3)$-DTS with sum-of-lengths
$\left(5+\frac{1}{91}\right)L^{2}+\left(1-\frac{15}{91}\right)L$
exists.
VI-D The Case of $M=4$
For $(L,4)$-DTSs, it is known that the trivial bound can be slightly strengthened by parity considerations to
$\max_{\ell\in[L]}d_{4}^{(\ell)}\geq\begin{cases}10L&\text{if }L\equiv 0\pmod{*}{2}\\ 10L+1&\text{if }L\equiv 1\pmod{*}{2}\end{cases}\text{.}$ (30)
Proposition 9.
The sum-of-lengths of an $(L,4)$-DTS satisfies
$\sum_{\ell=0}^{L-1}d_{4}^{(\ell)}\\ \geq\begin{dcases}9L^{2}+\frac{3}{2}L&\text{if }L\equiv 0\pmod{*}{2}\\ 9L^{2}+\frac{3}{2}L+\frac{1}{2}&\text{if }L\equiv 1\pmod{*}{2}\end{dcases}\text{.}$ (31)
Proof.
By the distinctness of positive differences, we have
$2\sum_{\ell=0}^{L-1}d_{4}^{(\ell)}=\sum_{\ell=0}^{L-1}\bigg{[}(d^{(\ell)}_{2}-d^{(\ell)}_{0})+(d^{(\ell)}_{4}-d^{(\ell)}_{2})+(d^{(\ell)}_{1}-d^{(\ell)}_{0})+\\ (d^{(\ell)}_{2}-d^{(\ell)}_{1})+(d^{(\ell)}_{3}-d^{(\ell)}_{2})+(d^{(\ell)}_{4}-d^{(\ell)}_{3})\bigg{]}\geq\sum_{u=1}^{6L}u$
which yields the result after integrality considerations. ∎
We conjecture that both (30) and (31) are simultaneously achievable for all $L$ excluding $L\in\{2,3,4\}$. We provide in the Appendix DTSs simultaneously achieving (30) and (31) for all $5\leq L\leq 10$ excluding $L=9$ for which a search is ongoing. The problem of simultaneously achieving (30) and (31) was considered implicitly by Laufer who solved it in the case of $L\in\{6,8\}$ in [46]. This was done because the structure implied by achieving (31) can be exploited to simplify the enumeration of solutions. A solution achieving both of (30) and (31) for $L=10$ was found by Khansefid et al. [47, 42, 48]. In the cases of $L\in\{5,7\}$, we found solutions with Shearer’s search programs [44, 45]. For $L\in\{2,3,4\}$, it is not only the case that one or both of (30) and (31) is not achievable, but also that we get a tradeoff between minimizing scope and minimizing sum-of-lengths when $L=4$. In this case, we provide a pair of Pareto optimal solutions. These smaller cases again can be handled by MILP or Shearer’s search programs [44, 45].
For all integers $i\geq 0$, let $\mathcal{Z}_{i}$ be a perfect $(L_{i},4)$-DTS with sum-of-lengths $S_{i}$ obtained by combining $\mathcal{Z}_{i-1}$ and $\mathcal{Z}_{0}$ via Theorem 5 where $\mathcal{Z}_{0}$ is a fixed perfect $(L_{0},4)$-DTS with sum-of-lengths $S_{0}$. Solving the linear recurrences for $L_{i}$ and $S_{i}$, we get
$L_{i}=\frac{1}{20}\left((20L_{0}+1)^{i+1}-1\right)$
and
$S_{i}=\frac{10S_{0}}{10L_{0}^{2}+L_{0}}L_{i}^{2}+\frac{S_{0}}{10L_{0}^{2}+L_{0}}L_{i}\text{.}$
[…]
$S_{i}=\left(9+\frac{6}{10L_{0}+1}\right)L_{i}^{2}+\left(\frac{3}{2}-\frac{6L_{0}}{10L_{0}+1}\right)L_{i}\text{.}$
Taking $\mathcal{Z}_{0}$ to be the perfect $(10,4)$-DTS achieving (31) in the Appendix, we get the following:
Proposition 10.
There are infinitely many values of $L$ for which a perfect $(L,4)$-DTS with sum-of-lengths
$\left(9+\frac{6}{101}\right)L^{2}+\left(\frac{3}{2}-\frac{60}{101}\right)L$
exists.
VI-E Implications for Higher-Order Staircase Codes
We now consider some implications of the results of this section for higher-order staircase codes:
VI-E1 Recovery of Tiled Diagonal Zipper Codes
Consider the higher-order staircase code defined by the $(2,S/L)$-nets from Example 1 or Example 2 and the scope and sum-of-lengths optimal $(L,1)$-DTS given by $d_{1}^{(\ell)}=L-\ell$ for $\ell\in[L]$. We get that
$\{d_{L},d_{L+1},\dots,d_{2L-1}\}$ =$=\{L(L-\ell)+\ell\mid\ell\in[L]\}$
=$=\{L(\ell+2)-(\ell+1)\mid\ell\in[L]\}$
which yields
$\pi^{\prime}_{L}=\pi^{\prime}_{L+1}=\dots=\pi^{\prime}_{2L-1}=\pi_{1}$
where $\pi_{1}(i,j)=(j,i)$. Our higher-order staircase code is then defined by the constraint that the rows of
$\big{(}B_{n-L^{2}}^{\mathsf{T}}\big{|}\cdots\big{|}B_{n-(3L-2)}^{\mathsf{T}}\big{|}B_{n-(2L-1)}^{\mathsf{T}}\big{|}B_{n-(L-1)}\big{|}\cdots\big{|}B_{n-1}\big{|}B_{n}\big{)}$
belong to some component code $\mathcal{C}$ for all $n\in L\mathbb{Z}$. This family is precisely the tiled diagonal zipper codes of [11].
VI-E2 Memory Reduction Limits
Tiled diagonal zipper codes are often colloquially described as providing encoding and decoding memory reductions by up to $1/2$ relative to classical staircase codes [2] which are the $L=1$ case for this family. The factor of $1/2$ arises as a certain reasonable limit representing the approximate memory reduction at high code rates. We can compute analogous limits for higher-order staircase codes. In particular, we divide the encoding memory (19) by the encoding memory in the $L=1$ case and take the limit as $S\to\infty$ with $S/L$ held constant so that $L\to\infty$ in each of the cases of $M\in\{1,2,3,4\}$. We will assume for each $M\in\{1,2,3,4\}$ the existence of infinitely many values of $L$ for which the respective sum-of-lengths lower bounds (25), (27), (29), and (31) are achieved. The resulting limits are $1/2$, $3/4$, $5/6$, and $9/11$ respectively for each $M\in\{1,2,3,4\}$. Note that our premise is only conjectured to be true in the cases of $M\in\{3,4\}$ but Propositions 8 and 10 allow us to prove limits of $5/6+1/546$ and $9/11+6/1111$ in those cases instead. We can also obtain decoding memory reduction limits via (20) similarly as $1/2$, $3/4$, $6/7$, and $10/12$ respectively for each $M\in\{1,2,3,4\}$.
VI-E3 Recovery of Self-Orthogonal Convolutional Codes
Consider a higher-order staircase code with $S=L$ and recall Definition 5 for $(M+1,1)$-nets. We see that the net permutations disappear from the construction. Consider further taking $r=1$, i.e., assuming single-parity-check constraints. The definition of higher-order staircase codes (Definition 10) still remains interesting in this case and leads to a family of rate $(L-1)/L$ codes. These codes are precisely a recursive rather than feedforward version of the self-orthogonal convolutional codes of Robinson and Bernstein [7]. In particular, the codes of [7] have a slightly increased rate of $L/(L+1)$ arising from an $(L,M)$-DTS achieved by omitting previous parities in the recursive encoding process. The reader is referred to [49] for an example of how the conversion to a feedforward variant can be done in the case of classical staircase codes. Such conversions typically lead to large error floors since parity symbols are left unprotected.
VII DTS Search Techniques
A highly-optimized variation on the DTS search algorithm of Koubi et al. [50] was implemented. Part of this variation is based on an adaptation of a standard programming trick in which a subset of $[N]$ is represented by a single $N$-bit integer. A critical, innermost loop step of the algorithm of Koubi et al. [50] and any similar stochastic local search algorithm is the insertion of a new mark into a partially constructed DTS. This requires that the distances created by the new mark are compared to all of the distances used by the partial DTS thus far, requiring up to $\mathcal{O}(LM^{3})$ operations. Under a C language model [51], this can be reduced to an essentially constant number of operations as will be seen shortly.
The resulting algorithm can also be used to make progress on the classical minimum scope DTS problem. As a certificate of its effectiveness, we exhibit a perfect (thus minimum scope) $(12,4)$-DTS and a minimum scope $(13,4)$-DTS settling the minimum scope DTS problems for these cases and improving on the previous best upper bounds of $122$ and $133$ given in [52]:
$\begin{array}[]{rrrrr}0&3&62&106&120\\ 0&11&66&86&119\\ 0&27&34&105&118\\ 0&18&56&99&116\\ 0&22&51&74&115\\ 0&10&42&77&114\\ 0&6&63&89&113\\ 0&2&47&87&112\\ 0&19&80&95&111\\ 0&21&70&100&109\\ 0&12&48&94&102\\ 0&28&96&97&101\\ \end{array}\quad\begin{array}[]{rrrrr}0&44&80&115&131\\ 0&33&42&123&130\\ 0&1&69&109&129\\ 0&25&73&84&127\\ 0&31&41&96&126\\ 0&26&64&78&125\\ 0&21&66&112&124\\ 0&8&83&100&122\\ 0&23&93&117&121\\ 0&15&82&101&119\\ 0&29&56&105&118\\ 0&6&63&113&116\\ 0&32&34&106&111\\ \end{array}$
VII-A C Language Model Speedup
We review key elements of the $C$ language model [51] to which the reader is referred for further details as needed:
  • An $N$-bit unsigned integer $\mathsf{n}$ is interpreted as an element of $\mathbb{Z}_{2^{N}}$ and its $i$th bit is the value of $\mathsf{b}_{i}\in\{0,1\}$ where $\mathsf{n}=\sum_{i=0}^{N-1}\mathsf{b}_{i}2^{i}$.
[…]
In what follows, all variables are $N$-bit unsigned integers which have a dual interpretation. When a variable is treated as a set, it represents the subset of $[N]$ whose elements are all $i\in[N]$ such that the $i$th bit of the variable is $1$. Otherwise, a variable simply represents its integer value. A partially constructed DTS is a set of rulers with differing numbers of elements such that all differences between elements within a ruler are distinct from each other and distinct from differences produced by other rulers. We are given a value $\mathsf{mark}$ which we wish to insert into a particular ruler $\mathsf{natRuler}$ of a partially constructed DTS, but only if it does not lead to repeated differences. We maintain:
  • the set of differences used by the partially constructed DTS thus far $\mathsf{usedDistances}$,
  • the largest mark of $\mathsf{natRuler}$ given by $\mathsf{largestMark}$, and
  • a redundant mirror image of $\mathsf{natRuler}$ given by $\mathsf{revRuler}$ where $\mathsf{revRuler}$ contains $i$ if and only if $\mathsf{natRuler}$ contains $\mathsf{largestMark}-i$.
If $\mathsf{mark}$ is successfully inserted into $\mathsf{natRuler}$, these three variables must also be updated accordingly.
Algorithm 1 Fast conditional mark insertion.
1:if $\mathsf{mark}>\mathsf{largestMark}$ then
2:     $\mathsf{leftDistances}\leftarrow\mathsf{revRuler}\ll(\mathsf{mark}-\mathsf{largestMark})$
3:     $\mathsf{rightDistances}\leftarrow 0$
4:else
5:     $\mathsf{leftDistances}\leftarrow\mathsf{revRuler}\gg(\mathsf{largestMark}-\mathsf{mark})$
6:     $\mathsf{rightDistances}\leftarrow\mathsf{natRuler}\gg\mathsf{mark}$
7:end if
8:$\mathsf{bisection}\leftarrow\mathsf{leftDistances}\mathbin{\&}\mathsf{rightDistances}$
9:$\mathsf{distances}\leftarrow\mathsf{leftDistances}\mathbin{|}\mathsf{rightDistances}$
10:$\mathsf{intersection}\leftarrow\mathsf{usedDistances}\mathbin{\&}\mathsf{distances}$
11:if $!(\mathsf{bisection}\mathbin{||}\mathsf{intersection})$ then
12:     $\mathsf{natRuler}\leftarrow\mathsf{natRuler}\mathbin{|}(1\ll\mathsf{mark})$
13:     if $\mathsf{mark}>\mathsf{largestMark}$ then
14:         $\mathsf{revRuler}\leftarrow\mathsf{leftDistances}\mathbin{|}1$
15:         $\mathsf{largestMark}\leftarrow\mathsf{mark}$
16:     else
17:         $\mathsf{revRuler}\leftarrow\mathsf{revRuler}\mathbin{|}(1\ll(\mathsf{largestMark}-\mathsf{mark}))$
18:     end if
19:     $\mathsf{usedDistances}\leftarrow\mathsf{usedDistances}\mathbin{|}\mathsf{distances}$
20:end if
We accomplish all of this with Algorithm 1. Here, $\mathsf{leftDistances}$ is assigned the distances between $\mathsf{mark}$ and all existing marks that are less than or equal to $\mathsf{mark}$. On the other hand, $\mathsf{rightDistances}$ is assigned the distances between $\mathsf{mark}$ and all existing marks that are greater than or equal to $\mathsf{mark}$. It can be shown that these two sets intersect if and only if the inserted $\mathsf{mark}$ is equal to an existing mark or bisects a pair of existing marks which leads to a repeated distance. Otherwise, the only possible repeated distances are common distances between $\mathsf{distances}$ and $\mathsf{usedDistances}$.
Finally, we must contend with the issue that in practice, we have $N=64$ in accordance with the $64$-bit data widths of modern processors. This apparently limits us to searching for DTSs with scope at most $64$. However, we can emulate a larger $N$ by implementing bitwise and shifting operations sequentially on $64$-bit chunks. This leads to increased complexity relative to what is presented as Algorithm 1, especially for the shifting operations which require carryovers. This also means that the number of operations is not strictly constant. However, there nonetheless remains a significant overall speedup relative to an array-based implementation performing mark-by-mark comparisons. An emulated $256$-bit implementation of Algorithm 1 can be found within our DTS search program sample implementation, available online [53].
VII-B Mark Statistics
The algorithm of Koubi et al. [50] involves a simple stochastic local search procedure in which a DTS is populated with randomly generated marks which are inserted if compatible as just described. Backtracking steps such as deleting an entire ruler or a mark within a ruler are used when no compatible marks are found after a specified number of attempts. These are accommodated by adding appropriate record-keeping procedures to Algorithm 1. However, the key novel aspect of the algorithm of Koubi et al. [50] is the modelling of the marks of a ruler in a DTS as arising from a simple Gaussian mixture: For each $i\in[M+1]$, the $i$th mark of any ruler in the DTS is sampled independently from a Gaussian distribution with mean $\mu_{i}$ and variance $\sigma^{2}_{i}$ (with rounding to the nearest integer). These parameters can be inferred in various ways described in [50] and a followup work [54].
We employ the following procedure for obtaining the mark distribution parameters $\mu_{i}$ and $\sigma^{2}_{i}$: Suppose that we seek DTSs with scope at most $T$. We run a version of our algorithm with a uniform mark distribution to find some number of DTSs with a larger admissible scope $\widetilde{T}$ with $\widetilde{T}>T$. If the resulting sample means and variances are $\widetilde{\mu}_{i}$ and $\widetilde{\sigma}^{2}_{i}$, we take $\mu_{i}=(T/\widetilde{T})\widetilde{\mu}_{i}$ and $\sigma_{i}^{2}=(T/\widetilde{T})^{2}\widetilde{\sigma}_{i}^{2}$. If we seek DTSs which further minimize sum-of-lengths, we add a third essential step: We generate a number of DTSs with scope at most $T$ and take $\mu_{i}$ and $\sigma^{2}_{i}$ to be the means and variances of a biased sample consisting only of those DTSs whose sum-of-lengths is below a specified threshold, e.g., below average. This results in an algorithm which is significantly more likely to produce a minimum sum-of-lengths DTS.
Aside from the key elements just described, there are some other minor differences from the algorithm of [50] which we will not bother describing; instead, we provide a sample implementation online [53].
VIII Future Work
Several directions for future work are immediately apparent. On the application side, while we have provided an essentially five-parameter ($L$, $M$, $S/L$, $C$, and $t$) family of codes, we have only evaluated performance for a few parameters with $t=1$ and with hard-decision decoding. Many other possibilities can be considered in future work such as $t>1$, soft-decision decoding, and any techniques that have been applied to product-like codes. Such investigations can be initiated by using and extending our simulators available online [21, 25] but can also be aided by the development of threshold and error floor analysis techniques along the lines of [3, 55, 26]. Furthermore, we may also consider the development of non-scattering and irregular-degree variants as done in [33] and [32] respectively.
On the mathematical side, it may be within reach to construct an infinite family of DTSs achieving the bounds (29) or (31) or to strengthen Propositions 8 and 10 by considering the many other construction methods for perfect DTSs provided in [41]. It may also be within reach to find an interesting general sum-of-lengths lower bound for arbitrary $M$. Lastly, the search for sum-of-lengths-optimal and scope-optimal DTSs can be continued.
[Explicit Difference Triangle Sets with Minimum Scope and Sum-of-Lengths]
-A $(L\geq 1,M=2)$
For $1\leq L\leq 7$, we have:
$\begin{array}[]{rrr}0&1&3\\ \end{array}\quad\begin{array}[]{rrr}0&2&7\\ 0&3&4\\ \end{array}\quad\begin{array}[]{rrr}0&3&10\\ 0&6&8\\ 0&4&5\\ \end{array}\quad\begin{array}[]{rrr}0&11&12\\ 0&4&10\\ 0&2&9\\ 0&3&8\\ \end{array}$
$\begin{array}[]{rrr}0&13&15\\ 0&6&14\\ 0&11&12\\ 0&3&10\\ 0&5&9\\ \end{array}\quad\begin{array}[]{rrr}0&3&19\\ 0&12&17\\ 0&6&15\\ 0&1&14\\ 0&7&11\\ 0&2&10\\ \end{array}\quad\begin{array}[]{rrr}0&2&22\\ 0&1&19\\ 0&6&17\\ 0&9&16\\ 0&12&15\\ 0&4&14\\ 0&5&13\\ \end{array}$
For all $L\geq 8$, we have:
-A1 $L=4m$ with $m>1$ due to Skolem [38]
$(0,4m-1,10m)$
$(0,2m-1,8m-1)$
$(0,1,5m+1)$
$(0,4m-2i,12m-i)\text{ for }i\in\{0,1,\dots,2m-1\}$
$(0,4m-1-2i,8m-1-i)\text{ for }i\in\{1,2,\dots,m-1\}$
$(0,2m-3-2i,7m-1-i)\text{ for }i\in\{0,1,\dots,m-3\}$
-A2 $L=4m+1$ with $m>1$ due to Skolem [38]
$(0,4m+1,10m+3)$
$(0,2m-1,8m+2)$
$(0,1,5m+3)$
$(0,4m-2i,12m+3-i)\text{ for }i\in\{0,1,\dots,2m-1\}$
$(0,4m+1-2i,8m+2-i)\text{ for }i\in\{1,2,\dots,m\}$
$(0,2m-1-2i,7m+2-i)\text{ for }i\in\{1,2,\dots,m-2\}$
-A3 $L=4m+2$ with $m>1$ due to O’Keefe [39]
$(0,4m+1,10m+4)$
$(0,2m+1,10m+5)$
$(0,4m+2,12m+7)$
$(0,1,11m+6)$
$(0,4m+2-2i,8m+4-i)\text{ for }i\in\{1,2,\dots,2m\}$
$(0,4m+1-2i,12m+6-i)\text{ for }i\in\{1,2,\dots,m-1\}$
$(0,2m+1-2i,11m+5-i)\text{ for }i\in\{1,2,\dots,m-1\}$
-A4 $L=4m+3$ with $m>1$ due to O’Keefe [39]
$(0,2m+3,7m+6)$
$(0,1,5m+5)$
$(0,2m+1,8m+6)$
$(0,4m+2,10m+8)$
$(0,4m+3,12m+10)$
$(0,4m+2-2i,12m+9-i)\text{ for }i\in\{1,2,\dots,2m\}$
$(0,4m+3-2i,8m+6-i)\text{ for }i\in\{1,2,\dots,m-1\}$
$(0,2m+1-2i,7m+6-i)\text{ for }i\in\{1,2,\dots,m-1\}$
-B $(1\leq L\leq 15,M=3)$
These are found with the methods described in Section VII.
$\begin{array}[]{rrrr}0&1&4&6\\ \end{array}\quad\begin{array}[]{rrrr}0&3&12&13\\ 0&5&7&11\\ \end{array}\quad\begin{array}[]{rrrr}0&3&15&19\\ 0&1&10&18\\ 0&2&7&13\\ \end{array}$
$\begin{array}[]{rrrr}0&6&14&24\\ 0&3&22&23\\ 0&9&16&21\\ 0&4&15&17\\ \end{array}\quad\begin{array}[]{rrrr}0&8&21&30\\ 0&14&26&29\\ 0&4&23&28\\ 0&7&25&27\\ 0&1&11&17\\ \end{array}\quad\begin{array}[]{rrrr}0&10&32&36\\ 0&8&24&35\\ 0&5&33&34\\ 0&12&25&31\\ 0&7&21&30\\ 0&2&17&20\\ \end{array}$
$\begin{array}[]{rrrr}0&14&30&42\\ 0&1&38&41\\ 0&8&32&39\\ 0&10&27&36\\ 0&13&33&35\\ 0&5&23&34\\ 0&4&19&25\\ \end{array}\quad\begin{array}[]{rrrr}0&2&43&48\\ 0&13&36&47\\ 0&15&33&45\\ 0&6&28&44\\ 0&7&39&42\\ 0&9&26&40\\ 0&8&27&37\\ 0&4&24&25\\ \end{array}\quad\begin{array}[]{rrrr}0&13&53&54\\ 0&4&49&52\\ 0&12&34&51\\ 0&14&43&50\\ 0&15&38&47\\ 0&16&35&46\\ 0&2&26&44\\ 0&6&27&37\\ 0&5&25&33\\ \end{array}$
$\begin{array}[]{rrrr}0&19&42&60\\ 0&15&43&59\\ 0&10&57&58\\ 0&17&49&56\\ 0&3&53&55\\ 0&20&46&54\\ 0&13&40&51\\ 0&9&31&45\\ 0&12&33&37\\ 0&5&29&35\\ \end{array}\quad\begin{array}[]{rrrr}0&10&53&66\\ 0&20&60&65\\ 0&2&63&64\\ 0&12&41&59\\ 0&16&44&58\\ 0&19&49&57\\ 0&7&33&55\\ 0&17&51&54\\ 0&21&46&52\\ 0&11&35&50\\ 0&9&32&36\\ \end{array}\quad\begin{array}[]{rrrr}0&14&59&72\\ 0&16&49&71\\ 0&5&67&70\\ 0&21&52&69\\ 0&24&53&68\\ 0&9&60&66\\ 0&23&63&64\\ 0&11&43&61\\ 0&20&46&56\\ 0&7&35&54\\ 0&4&34&42\\ 0&12&37&39\\ \end{array}$
$\begin{array}[]{rrrr}0&17&66&78\\ 0&22&62&77\\ 0&2&73&76\\ 0&10&68&75\\ 0&20&47&72\\ 0&24&54&70\\ 0&21&56&69\\ 0&23&59&67\\ 0&19&60&64\\ 0&26&57&63\\ 0&11&39&53\\ 0&18&50&51\\ 0&9&38&43\\ \end{array}\quad\begin{array}[]{rrrr}0&6&77&84\\ 0&18&57&83\\ 0&8&80&82\\ 0&23&53&81\\ 0&12&55&79\\ 0&3&63&76\\ 0&27&64&75\\ 0&14&45&70\\ 0&22&54&69\\ 0&19&52&68\\ 0&20&62&66\\ 0&17&51&61\\ 0&21&50&59\\ 0&5&40&41\\ \end{array}\quad\begin{array}[]{rrrr}0&5&76&90\\ 0&1&82&89\\ 0&21&58&87\\ 0&19&59&86\\ 0&4&64&84\\ 0&28&72&83\\ 0&18&63&79\\ 0&24&70&78\\ 0&15&51&77\\ 0&23&73&75\\ 0&25&57&74\\ 0&13&47&69\\ 0&30&65&68\\ 0&12&43&53\\ 0&6&39&48\\ \end{array}$
-C $(1\leq L\leq 10,L\neq 9,M=4)$
The cases of $L\in\{6,8\}$ are due to Laufer [46, 42, 48], the case of $L=10$ is due to Khansefid et al. [47, 42, 48], and the cases of $L\in\{5,7\}$ were found with Shearer’s programs [44, 45].
$\begin{array}[]{rrrrr}0&1&4&9&11\\ \end{array}\quad\begin{array}[]{rrrrr}0&2&9&21&22\\ 0&4&10&15&18\\ \end{array}$
$\begin{array}[]{rrrrr}0&2&10&19&32\\ 0&3&15&26&31\\ 0&1&7&21&25\\ \end{array}\quad\begin{array}[]{rrrrr}0&4&16&34&41\\ 0&13&23&32&40\\ 0&3&24&38&39\\ 0&5&11&31&33\\ \end{array}$
$\begin{array}[]{rrrrr}0&5&19&40&42\\ 0&7&15&33&39\\ 0&9&22&34&38\\ 0&1&11&28&31\\ \end{array}\quad\begin{array}[]{rrrrr}0&6&20&48&51\\ 0&9&21&46&50\\ 0&13&23&47&49\\ 0&5&16&35&43\\ 0&1&18&33&40\\ \end{array}$
$\begin{array}[]{rrrrr}0&14&26&51&60\\ 0&4&28&44&59\\ 0&10&23&52&58\\ 0&1&21&54&57\\ 0&7&18&45&50\\ 0&2&19&41&49\\ \end{array}\quad\begin{array}[]{rrrrr}0&8&28&67&71\\ 0&10&33&57&70\\ 0&5&34&55&69\\ 0&12&27&65&68\\ 0&1&26&45&62\\ 0&7&18&49&58\\ 0&6&22&52&54\\ \end{array}$
$\begin{array}[]{rrrrr}0&19&34&73&80\\ 0&8&35&63&79\\ 0&12&33&74&78\\ 0&13&30&72&77\\ 0&11&36&67&76\\ 0&18&32&69&75\\ 0&2&22&60&70\\ 0&1&24&50&53\\ \end{array}\quad\begin{array}[]{rrrrr}0&1&45&98&100\\ 0&9&36&77&96\\ 0&14&37&88&95\\ 0&10&35&83&94\\ 0&15&46&76&93\\ 0&12&40&79&92\\ 0&22&42&85&91\\ 0&8&34&72&90\\ 0&3&32&65&89\\ 0&5&21&71&75\\ \end{array}$